### v1

# Positroid envelopes and graphic positroids

Abstract.
Positroids are matroids realizable by real matrices with all nonnegative maximal minors. They partition the ordered matroids into equivalence classes, called positroid envelope classes, by their Grassmann necklaces. We give an explicit graph construction that shows that every positroid envelope class contains a graphic matroid. We show that the following classes of positroids are equivalent: graphic, binary, and regular, and that a graphic positroid is the unique matroid in its positroid envelope class. Finally, we show that every graphic positroid has an oriented graph representable by a signed incidence matrix with all nonnegative minors.
1. Introduction
The Grassmannian $\operatorname{Gr}(k,\mathbb{F}^{n})$ admits a stratification by the realization spaces of the rank-$k$ ordered matroids on $n$ elements. The study of this matroid stratification of the Grassmannian has proven difficult as, due to Mnëv’s Universality Theorem, a matroid stratum can be as pathological as any algebraic variety. The totally nonnegative Grassmannian $\operatorname{Gr}^{\geq 0}(k,\mathbb{R}^{n})$, consisting of the points in $\operatorname{Gr}(k,\mathbb{R}^{n})$ that can be realized by matrices with all nonnegative maximal minors, has a stratification due to Postnikov into positroid cells that are each homeomorphic to an open ball [14].
Positroids are in bijection with a number of interesting combinatorial objects [14], and have connections to scattering amplitudes in $\mathcal{N}=4$ supersymmetric Yang-Mills theory [3], tilings of the amplituhedron [12], and cluster algebras [15]. Postnikov’s positroid stratification of the nonnegative Grassmannian yields a partition of the ordered matroids into classes with positroid representatives, which are called the positroid envelopes [5]. We study the interplay between the family of graphic matroids and these positroid envelopes. We show that every positroid is the envelope of a graphic matroid. We show that the following are equivalent: $P$ is a graphic positroid; $P$ is binary; $P$ is regular; $P$ is the unique matroid contained in its positroid envelope class.
The rest of this paper is organized as follows. In Section 2, we provide the necessary definitions and tools. In Section 3, we prove that every positroid envelope class contains at least one graphic matroid, by explicitly constructing a planar graph whose cycle matroid lies in the desired positroid envelope. In Section 4, we characterize the graphic positroids and their envelope classes. In Section 5, we strengthen the connection between the graphic positroids and their graphs, by showing that every graphic positroid $P$ can be realized by a matrix with all nonnegative maximal minors that is the signed incidence matrix of a graph that represents $P$.
2. Basic tools and definitions
In this section, we establish basic definitions and tools that are used throughout the paper. In 2.1, we establish the basics for matroids. In 2.2, we give necessary background on positroids. In 2.3, we describe some combinatorial objects that are in bijection to the positroids: decorated permutations and Grassmann necklaces. These are used in the proofs of main results. In 2.4, we give the basic definitions of positroid envelopes and varieties.
2.1. Matroids
A matroid $M=(E,\mathcal{B})$ consists of a finite ground-set $E$ and a non-empty collection of subsets $\mathcal{B}\subset 2^{E}$, called bases, that satisfy the following basis exchange property:
  • if $B_{1},B_{2}\in\mathcal{B}$ and $x\in B_{1}\setminus B_{2}$, then there exists $y\in B_{2}\setminus B_{1}$ such that $(B_{1}-x)\cup\{y\}\in\mathcal{B}$.
The collection of independent sets of $M$ is $\mathcal{I}(M):=\{I\subseteq B:B\in\mathcal{B}(M)\}$, and the dependent sets of $M$ are all subsets of $E(M)$ that are not independent. Minimal dependent sets of $M$ are called circuits, and a circuit consisting of exactly one element is called a loop. The dual of a matroid $M$ is the matroid $M^{*}$ whose ground-set is $E(M^{*}):=E(M)$ and whose bases are $\mathcal{B}(M^{*}):=\{E(M)-B:B\in\mathcal{B}(M)\}$. A cocircuit of $M$ is a circuit in $M^{*}$, and a coloop of $M$ is a loop in $M^{*}$. The rank function of a matroid $M$ is a map $r_{M}:2^{E(M)}\to\mathbb{Z}_{\geq 0}$ given by, for all $X\subseteq E(M)$, $r_{M}(X):=\max\{\lvert I\rvert:I\subseteq X,I\in\mathcal{I}(M)\}$. We call $X\subseteq E(M)$ a flat of $M$ if $\forall y\in E(M)-X,r_{M}(X\cup y)>r_{M}(X)$. A flat $X$ of $M$ is called a hyperplane if $r_{M}(X)=r_{M}(M)-1$. We call $X\subset E(M)$ a circuit-hyperplane if $X$ is both a circuit and a hyperplane of $M$. The closure operator of a matroid $M$ is a function $\operatorname{cl}:2^{E}\to 2^{E}$ given by, for all $X\subseteq E(M)$, $\operatorname{cl}(X):=\{e\in E(M):r_{M}(X\cup e)=r_{M}(X)\}$. A matroid isomorphism between two matroids $M$ and $N$ is a basis-preserving bijection $E(M)\leftrightarrow E(N)$.
Let $M$ be a matroid and $x\in E(M)$. The matroid obtained by deleting $x$ from $M$, $M\setminus x$, has ground-set $E(M\setminus x):=E(M)\setminus\{x\}$ and independent sets $\mathcal{I}(M\setminus x):=\{I:x\notin I\in\mathcal{I}(M)\}$. The matroid obtained by contracting $x$ in $M$, $M/x$, has ground-set $E(M/x):=E(M)\setminus\{x\}$ and bases
$\mathcal{B}(M/x):=\begin{cases}\{B:B\in\mathcal{B}(M)\},&\text{if }x\text{ is a loop}\\ \{B\setminus\{x\}:x\in B\in\mathcal{B}(M)\},&\text{otherwise.}\end{cases}$
We use the corresponding notation $(M/X)\setminus Y$, for $X,Y\subseteq E(M)$ and $X\cap Y=\emptyset$, to mean the contraction of the set of elements $X$ and deletion of the set of elements $Y$. Deletion and contraction are dual to each other, in the sense that $M\setminus x=(M^{*}/x)^{*}$ and $M/x=(M^{*}\setminus x)^{*}$. We say that a matroid $N$ is a minor of $M$ if it can be obtained from $M$ by a sequence of deletions and contractions. We say that a matroid $M$ is $N$-free if $N$ is not a minor of $M$.
Let $A$ be a matrix with entries over a field $\mathbb{F}$. We can obtain a matroid, $M(A)$, from $A$ by taking the ground-set to be the column vectors of $A$, and the bases to be the collection of maximal linearly independent subsets of the column vectors of $A$. We say that $A$ is a matrix that realizes the matroid $M(A)$. If there is an isomorphism between a matroid $M$ and $M(A)$, then we also say that $A$ realizes $M$. A matroid is called $\mathbb{F}$-linear, or $\mathbb{F}$-realizable, if it can be realized by a matrix with entries over $\mathbb{F}$. If a matroid $M$ is $\mathbb{F}$-linear, then so to is $M^{*}$. A matroid is called binary or ternary if it is $\mathbb{F}_{2}$ or $\mathbb{F}_{3}$-linear, respectively. A matroid $M$ is regular if it is $\mathbb{F}$-linear for all fields $\mathbb{F}$. For any field $\mathbb{F}$, the class of $\mathbb{F}$-linear matroids is minor-closed, thus can be characterized by a (possibly infinite) set of forbidden minors. For example, binary matroids have a simple characterization.
Theorem 2.1 (Theorem 6.5.4 in [11]).
A matroid is binary if and only if it has no minor isomorphic to $U^{2}_{4}$.
A graph $G=(V,E)$ on $n$ edges consists of a vertex set $V$ and an edge set $E$. Edges have either two vertex endpoints or one, in which case they are called loops. We allow multiple edges to have the same endpoints. We always take $E$ to have a total order. A cycle in $G$ is a set of vertices $v_{1},\dots,v_{k}\in V(G)$ and edges $e_{1},\dots,e_{k}\in E(G)$ such that each $e_{i}$ has endpoints $v_{i}$ and $v_{i+1}$ with indices taken modulo $k$. Given a graph $G$, the collection of maximal forests of $G$ give the bases of a matroid on $E$, which we denote as $M(G)$. We say that $G$ represents a matroid $M$ if there is an isomorphism between $M$ and $M(G)$. A graphic matroid $M$ is a matroid that can be represented by some graph. Every graphic matroid is regular. Graphic matroids have the following forbidden minor characterization.
Theorem 2.2 (Theorem 10.3.1 in [11]).
A matroid is graphic if and only if it has no minor isomorphic to any of the matroids $U^{2}_{4}$, $F_{7}$, $F^{*}_{7}$, $M^{*}(K_{5}),$ or $M^{*}(K_{3,3})$.
2.1.1. Connectivity, direct sums and $2$-sums
Let $M$ be a matroid and $X\subseteq E(M)$. The connectivity function of $M$, $\lambda_{M}$, is defined as
$\lambda_{M}(X):=r_{M}(X)+r_{M}(E\setminus X)-r_{M}(M).$
A $k$-separation of $M$ is a pair $(X,E\setminus X)$ for which $\min\{\lvert X\rvert,\lvert E\setminus X\rvert\}\geq k$ and $\lambda_{M}(X)<k$. For $n\geq 2$, a matroid $M$ is $n$-connected if, for all $k\in[n-1]$, $M$ has no $k$-separations. We call a matroid $M$ disconnected if it is not $2$-connected.
Let $M$ and $N$ be matroids on disjoint sets. The direct sum of $M$ and $N$, $M\oplus N$, is the matroid whose ground-set is $E(M\oplus N):=E(M)\cup E(N)$ and whose bases are $\mathcal{B}(M\oplus N):=\{B\cup B^{\prime}:B\in\mathcal{B}(M),B^{\prime}\in\mathcal{B}(N)\}$.
A matroid is disconnected if and only if it is the direct sum of two non-empty matroids. Every matroid $M$ decomposes under direct sums into a unique collection of non-empty $2$-connected matroids which we call the $2$-connected components (or connected components) of $M$. We denote the number of $2$-connected components of a matroid $M$ by $c(M)$.
Let $M$ and $N$ be matroids, such that
  • (i)
    $\lvert E(M)\rvert,\lvert E(N)\rvert\geq 2$,
  • (ii)
    $E(M)\cap E(N)=\{e\}$, and
  • (iii)
    neither $(e,E(M)\setminus\{e\})$ nor $(e,E(N)\setminus\{e\})$ is a $1$-separator of $M$ or $N$ respectively.
Then the $2$-sum of $M$ and $N$, $M\oplus_{2}N$, is the matroid on $E(M\oplus_{2}N)=(E(M)\cup E(N))\setminus\{e\}$ whose circuits are
$\mathcal{C}(M\setminus e)$ $\cup\mathcal{C}(N\setminus e)$
$\cup\{(C_{1}\cup C_{2})\setminus\{e\}:C_{1}\in\mathcal{C}(M),C_{2}\in\mathcal{C}(N),e\in C_{1}\cap C_{2}\}.$
We call a $2$-sum, $M=M_{1}\oplus_{2}M_{2}$, non-trivial if $|E(M_{1})|,|E(M_{2})|>2$.
From a matroid $M=M_{1}\oplus_{2}M_{2}$, we can obtain the matroids $M_{1}$ and $M_{2}$ as minors by the following result.
Proposition 2.3 (Proposition 7.1.21 in [11]).
Both $M$ and $N$ are isomorphic to minors of $M\oplus_{2}N$. In particular, if $p^{\prime}$ is an element of $E(N)\setminus p$ that is in the same component of $N$ as $p$, then there is a minor of $M\oplus_{2}N$ from which $M$ can be obtained by relabelling $p^{\prime}$ by $p$. Moreover, if $N$ has at least three elements, then $M$ isomorphic to a proper minor of $M\oplus_{2}N$.
2.1.2. Tree decompositions
Every matroid $M$ decomposes under direct sums and $2$-sums into a collection of maximal non-empty $3$-connected matroids which we call the $3$-connected components of $M$. Such a decomposition can be represented by a vertex- and edge-labelled tree.
Definition 2.4 ([11]).
A matroid-labelled tree is a tree $T$ with vertex set $\{M_{1},M_{2},\ldots,M_{k}\}$ for some positive integer $k$ such that
  • (i)
    each $M_{i}$ is a matroid;
  • (ii)
    if $M_{i}$ and $M_{j}$ are joined by an edge $e$ of $T$, then $E(M_{i})\cap E(M_{j})=\{e\}$, where $e$ is not a $1$-seperator of $M_{i}$ or $M_{j}$; and
  • (iii)
    if $M_{i}$ and $M_{j}$ are non-adjacent, then $E(M_{i})\cap E(M_{j})=\emptyset$.
Let $T$ be a matroid-labelled tree and let $e\in E(T)$ with endpoints $M_{i}$ and $M_{j}$. Then the tree $T/e$, obtained from $T$ by contracting the edge $e$ and labelling with $M_{i}\oplus_{2}M_{j}$ the new vertex that arises from identifying $M_{i}$ and $M_{j}$, is a matroid-labelled tree.
Let $e$ be an edge in some matroid-labelled tree $T$ with endpoints $M_{i}$ and $M_{j}$. A relabelling move of $e$ in $T$ consists of relabelling $e\in E(T),E(M_{i}),$ and $E(M_{j})$ by some $f$. We say that two matroid-labelled trees are equal to within relabelling of their edges if one can be obtained from the other by a sequence of relabelling moves.
Definition 2.5 ([11]).
A tree decomposition of a $2$-connected matroid $M$ is a matroid-labelled tree $T$ such that if $V(T)=\{M_{1},M_{2},\ldots,M_{k}\}$ and $E(T)=\{e_{1},e_{2},\ldots,e_{k-1}\}$, then
  • (i)
    $E(M)=(E(M_{1})\cup E(M_{2})\cup\cdots\cup E(M_{k}))\setminus\{e_{1},e_{2},\ldots,e_{k-1}\}$;
  • (ii)
    $\lvert E(M_{i})\rvert\geq 3$ for all $i$ unless $\lvert E(M)\rvert<3$, in which case $k=1$ and $M_{1}=M$; and
  • (iii)
    $M$ is the matroid that labels the single vertex of $T/\{e_{1},\ldots,e_{k-1}\}$.
Theorem 2.6 (Theorem 8.3.10 in [11]).
Let $M$ be a $2$-connected matroid. Then $M$ has a tree decomposition $T$ in which every vertex label is $3$-connected, a circuit, or a cocircuit, and there are no two adjacent vertices that are both labelled by circuits or both labelled by cocircuits. Moreover, $T$ is unique to within relabelling of its edges.
For a given $2$-connected matroid $M$, we call the tree decomposition given by Theorem 2.6 the canonical tree decomposition for $M$. The reader is directed to Section 8.3 in [11] for more details on matroid-labelled trees and tree decompositions of $2$-connected matroids.
2.1.3. Series-parallel matroids
Let $M$ and $N$ be matroids such that $E(M)\cap E(N)=\{e\}$ and $e$ is not a $1$-separator of $M$ or $N$. The series-connection of $M$ and $N$ is the matroid $S(M,N)$ whose ground-set is $E(S(M,N)):=E(M)\cup E(N)$ and whose set of circuits is
$\mathcal{C}(S(M,N)):=\mathcal{C}(M\setminus e)\cup\mathcal{C}(N\setminus e)\cup\{C_{1}\cup C_{2}:e$ $\in C_{1}\in\mathcal{C}(M),$
$e$ $\in C_{2}\in\mathcal{C}(N)\}.$
The parallel connection of $M$ and $N$ is the matroid $P(M,N)$ whose ground-set is $E(P(M,N)):=E(M)\cup E(N)$ and whose set of circuits is
$\mathcal{C}(P(M,N)):=\mathcal{C}(M)\cup\mathcal{C}(N)\cup\{(C_{1}\cup C_{2})\setminus\{e\}):e$ $\in C_{1}\in\mathcal{C}(M),$
$e$ $\in C_{2}\in\mathcal{C}(N)\}.$
If $e$ is a loop of $M$ then
$P(M,N)$ $=P(N,M)=M\oplus(N/e),\text{ and}$
$S(M,N)$ $=S(N,M)=(M/e)\oplus N.$
If $e$ is a coloop of $M$ then
$P(M,N)$ $=P(N,M)=(M\setminus e)\oplus N,\text{ and}$
$S(M,N)$ $=S(N,M)=M\oplus(N\setminus e).$
We say that $N$ is a series extension of $M$ if $N\cong S(M,U^{1}_{2})$ and we say that $N$ is a parallel extension of $M$ if $N\cong P(M,U^{1}_{2})$. A matroid $M$ is a series-parallel matroid if $M$ can be obtained from $U^{0}_{1}$ or $U^{1}_{1}$ by a sequence of series and parallel extensions.
2.1.4. Weak maps
Let $M$ and $N$ be matroids. A weak map from $M$ to $N$ is a bijection $\varphi:E(M)\to E(N)$ such that $\forall I\in\mathcal{I}(N)$, $\varphi^{-1}[I]\in\mathcal{I}(M)$. We say that a weak map $\varphi:E(M)\to E(N)$ is rank-preserving if $r_{M}(M)=r_{N}(N)$. A rank-preserving weak map between two matroids implies the following connectivity result.
Proposition 2.7.
Let $M$ and $N$ be matroids such that $\mathbb{1}:M\to N$ is a rank-preserving weak map. If $N$ is $n$-connected, then $M$ is $n$-connected.
Proof.
If $\mathbb{1}:M\to N$ is a weak map, then by Proposition 9.1.2(e) in [17], for every subset $X\subseteq E$, we have $r_{M}(X)\geq r_{N}(X)$. If $\mathbb{1}:M\to N$ is rank-preserving, it follows that $\lambda_{M}(X)\geq\lambda_{N}(X)$ for all $X\subseteq E$. ∎
A rank-$k$ matroid $M$ on $n$ elements can be represented by the matroid base polytope $\Gamma_{M}$, which is a subset of the hypersimplex $\Delta_{n,k}$. We can obtain the number of connected components of $M$ from $n$ and the dimension of $\Gamma_{M}$.
Definition 2.8.
Let $M$ be a matroid on $\{i_{1}<i_{2}<\cdots<i_{n}\}$. For $I\in\mathcal{I}(M)$, let $e_{I}=\sum_{i_{j}\in I}e_{j}$, where $\{e_{1},e_{2},\ldots,e_{n}\}\subset\mathbb{R}^{n}$ are the standard basis vectors. The matroid base polytope of $M$ is
$\Gamma_{M}=\operatorname{convex}\{e_{B}:B\in\mathcal{B}(M)\}\subset\mathbb{R}^{n}.$
Corollary 2.9 (Corollary 4.7 in [4]).
Let $M$ be a matroid on the set $\{i_{1}<\ldots<i_{n}\}$ and $c(M)$ the number of connected components of $M$. Then the dimension of the basis matroid polytope associated with $M$ is given by the formula
$\dim\Gamma_{M}=n-c(M).$
We can use a rank-preserving weak map $\mathbb{1}:M\to N$ to bound the number of connected components of $M$ above by the number of connected components of $N$.
Corollary 2.10.
Let $M$ and $N$ be matroids, such that $\mathbb{1}:M\to N$ is a rank preserving weak map. Then, $c(M)\leq c(N)$.
Proof.
The rank-preserving weak map $\mathbb{1}:M\to N$ implies that $\mathcal{B}(N)$ is contained in $\mathcal{B}(M)$. Therefore, the matroid base polytope $\Gamma_{N}$ is contained in $\Gamma_{M}$. Thus, the dimension of $\Gamma_{N}$ is at most the dimension of $\Gamma_{M}$. Then, by Corollary 2.9,
$|E(N)|-c(N)\leq|E(M)|-c(M).$
It follows that $c(M)\leq c(N)$. ∎
Rank-preserving weak maps interact nicely with matroid minors. Let $M$ be a matroid $X,Y$ disjoint subsets of $E(M)$. The matroid minor $N=(M/X)\setminus Y$ is called properly expressed if it is obtained from $M$ without contracting any loops of deleting any coloops (Definition 5.6 in [6]).
Proposition 2.11 (Proposition 5.7 in [6]).
Any minor can be properly expressed.
We can use a rank-preserving weak map $\mathbb{1}:M\to N$ and a properly expressed minor $N^{\prime}$ of $N$ to find a minor $M^{\prime}$ of $M$ for which $\mathbb{1}:M^{\prime}\to N^{\prime}$ is a rank-preserving weak map.
Theorem 2.12 (Theorem 5.8 in [6]).
Let $M$ and $N$ be matroids such that $\mathbb{1}:M\to N$ is a rank-preserving weak map. If $X,Y\subseteq E(M)=E(N)$ and $N^{\prime}=(N/X)\setminus Y$ is a properly expressed minor of $N$, then $M^{\prime}=(M/X)\setminus Y$ is a minor of $M$ such that $\mathbb{1}:M^{\prime}\to N^{\prime}$ is a rank-preserving weak map.
The uniform matroid $U^{k}_{n}$ is the matroid on $[n]$ whose bases are all $k$-element subsets of $[n]$. For any matroid $N$ on $X$ of rank-$r$, where $r\leq k$, we have $\mathcal{B}(N)\subseteq\mathcal{B}(U^{k}_{n})$. This implies the following result.
Corollary 2.13 (Corollary 5.9 in [6]).
Let $M$ and $N$ be matroids such that $\mathbb{1}:M\to N$ is a rank-preserving weak map. If $M$ is $U^{k}_{n}$-free then $N$ is $U^{k}_{n}$-free.
By Theorem 2.1, a matroid is binary if and only if it is $U^{2}_{4}$-free, so we have the following.
Theorem 2.14 (Theorem 6.5 in [6]).
Let $M$ and $N$ be matroid such that $\mathbb{1}:M\to N$ is a rank-preserving weak map. If $M$ in binary then so is $N$.
A rank-preserving weak map $\mathbb{1}:M\to N$ is called nontrivial if $M\neq N$. If $M$ is binary and $N$ is $2$-connected, then there is no nontrivial rank-preserving weak map $\mathbb{1}:M\to N$.
Theorem 2.15 (Theorem 6.10 in [6]).
Let $M$ and $N$ be matroids such that $\mathbb{1}:M\to N$ is a rank-preserving weak map. If $M$ is binary, then $N$ is disconnected.
2.2. Positroids
We now consider the main object of our study, a class of ordered matroids called positroids. An ordered matroid $M$ is a matroid whose ground-set $E(M)$ has a total ordering. We say that a matroid isomorphism $\phi:M\to N$ is order preserving if both $M$ and $N$ are ordered matroids, and $\forall\{x<y\}\subseteq E(M),\{\phi(x)<\phi(y)\}\subseteq E(N)$. Recall that the maximal minors of a matrix $A$ is the collection of determinants of all maximal square submatrices of $A$. If $A$ is a square matrix, then the unique maximal minor is the determinant of $A$.
Definition 2.16.
A positroid is an $\mathbb{R}$-linear ordered matroid that can be realized by a matrix $A$ over $\mathbb{R}$ with all nonnegative maximal minors.
The following proposition due to Ardila, Rincón, and Williams shows that positroids are closed under duality and taking minors.
Proposition 2.17 (Proposition 3.5 in [1]).
Let $M$ be a positroid on $\{i_{1}<\ldots<i_{n}\}$. Then $M^{*}$ is a positroid on $\{i_{1}<\ldots<i_{n}\}$. Furthermore, for any subset $S$ of $\{i_{1}<\ldots<i_{n}\}$, the deletion $M\setminus S$ and the contraction $M/S$ are both positroids on $E(M)\setminus S$. Here the total order on $E(M)-S$ is the one inherited from $\{i_{1}<\ldots<i_{n}\}$.
Let $X$ and $Y$ be subsets of a totally ordered set $Z=\{i_{1}<i_{2}<\cdots<i_{n}\}$. For a fixed $i_{k}\in Z$, we define a new total order $\leq_{i_{k}}$ on $Z$ given by
$i_{k}<i_{k+1}<\cdots i_{n}<i_{1}<i_{2}<\cdots i_{k-1}.$
We say that $X$ and $Y$ are crossing subsets of $Z$ if there exist $w,y\in X$ and $x,z\in Y$ such that $w<_{w}x<_{w}y<_{w}z$. If $X$ and $Y$ are not crossing subsets of $Z$, then we call them non-crossing subsets of $Z$. The following characterization of positroids, using non-crossing subsets, is due to Ardila, Rincón, and Williams, building on prior work of da Silva.
Theorem 2.18 (Theorem 5.1,5.2 in [2]; Chapter 4, Theorem 1.1 in [13]).
An ordered matroid $M$ on the set totally ordered set $E(M)$ is a positroid if and only if for any circuit $C$ and any cocircuit $C^{*}$ satisfying $C\cap C^{*}=\emptyset$, the sets $C$ and $C^{*}$ are non-crossing subsets of $E(M)$.
Let $X$ be a totally ordered set and $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{t}\}$ be a partition of $X$. Then $\mathcal{S}$ is a non-crossing partition if for every pair $i,j\in[t]$, where $i\neq j$, $S_{i}$ and $S_{j}$ are non-crossing subsets of $X$. Ardila, Rincón, and Williams presented the following characterization of disconnected positroids using non-crossing partitions.
Theorem 2.19 (Theorem 7.6 in [1]).
Let $M$ be a positroid on $X$ and let $S_{1},S_{2},\ldots,S_{t}$ be the ground-sets of the connected components of $M$. Then $\{S_{1},\ldots,S_{t}\}$ is a non-crossing partition of $X$. Conversely, if $S_{1},S_{2},\ldots,S_{t}$ form a non-crossing partition of $X$ and $M_{1},M_{2},\ldots,M_{t}$ are connected positroids on $S_{1},S_{2},\ldots,S_{t}$, respectively, then $M_{1}\oplus\cdots\oplus M_{t}$ is a positroid.
2.2.1. Dihedral action
Let $Z=\{i_{1}<i_{2}<\cdots<i_{n}\}$. Let $r$ and $s$ be the bijections on $Z$ given by $r(i_{k})=r(i_{k+1})$ and $s(i_{k})=s(i_{n-k+1})$, where the indices are taken modulo $n$. Take $D_{Z}$ to be the (dihedral) group of bijections generated by $r$ and $s$ under function composition. We define the group action of $D_{Z}$ on the powerset $2^{Z}$ as follows, for all $\omega\in D_{Z}$ and any subset $X\subseteq Z$,
$\omega\cdot X=\{\omega(x):x\in X\}.$
We define the group action of $D_{Z}$ on the set {ordered matroids on $Z$} as follows, for all $\omega\in D_{Z}$ and any ordered matroid $M$ on $Z$, the ordered matroid $\omega\cdot M$ has ground-set $\omega\cdot Z=Z$ and set of bases
$\mathcal{B}(\omega\cdot M):=\{\omega[B]:B\in\mathcal{B}(M)\}.$
Non-crossing subsets, and therefore positroids, are preserved under dihedral action. The closure of positroids under dihedral action is a known folklore result. We include a proof for completeness.
Proposition 2.20.
Let $M$ be a positroid on $Z=\{i_{1}<\ldots<i_{n}\}$. Then, for any $\omega\in D_{Z}$, $\omega\cdot M$ is also a positroid.
Proof.
The preservation of positroids under $r$ is shown in [1]. Let $X,Y\subseteq Z$. Suppose that $s\cdot X$ and $s\cdot Y$ are crossing subsets of $s\cdot Z=Z$. Then, there exist $w,y\in s\cdot X$ and $x,z\in s\cdot Y$ such that $w<_{w}x<_{w}y<_{w}z$. Hence, $s(w),s(y)\in X$ and $s(x),s(z)\in Y$. Furthermore
$s(z)<_{s(z)}s(y)<_{s(z)}s(x)<_{s(z)}s(w).$
Therefore, $X$ and $Y$ are crossing subsets of $Z$. This completes the proof that the dihedral action of $D_{Z}$ preserves positroids. ∎
2.2.2. Series-parallel connections and 2-sums of positroids
The following characterization of non-crossing subsets is useful when we consider $2$-sums and series-parallel connections of positroids. Instead of just assuming that the elements of a matroid are ordered, it is convenient to assume that they are a subset of a densely ordered set. Then, it is always possible to add an element to the set anywhere in the order.
Observation 2.21.
Let $X$ and $Y$ be subsets of a densely totally ordered set $Z$ such that $\lvert X\rvert\geq 2$, $\lvert Y\rvert\geq 2$, and $X\cap Y=\{e\}$. Then $X$ and $Y$ are non-crossing subsets of $Z$ if and only if
  • (i)
    $\forall x\in X$, $\forall y\in Y$, $x\leq_{x}e\leq_{x}y$; or
  • (ii)
    $\forall x\in X$, $\forall y\in Y$, $y\leq_{y}e\leq_{y}x$.
Proof.
Without loss of generality, suppose that $\forall x\in X$, $\forall y\in Y$, $x\leq_{x}e\leq_{x}y$. For distinct $a,b\in X$ and $c,d\in Y$ we have either $a<_{a}b\leq_{a}c,d$ or $b<_{b}a\leq_{b}c,d$. Therefore, $X$ and $Y$ are non-crossing.
Now, suppose that $X$ and $Y$ are non-crossing subsets of $Z$ and that neither (i) nor (ii) hold. Then there exist $x\in X-e$ and $y\in Y-e$ such that $x<_{x}y<_{x}e$. Let
$x^{\prime}=\max_{\leq_{e}}\{x\in X-e:x<_{x}y<_{x}e\}.$
Then there does not exist any $x^{\prime\prime}$ such that $x^{\prime}<_{x}x^{\prime\prime}<_{x}y$, and, since $X$ and $Y$ are non-crossing, there does not exist any $y^{\prime}\in Y$ such that $e<_{x}y^{\prime}<_{x}x$ and there does not exist any $x^{\prime\prime}$ such that $y<_{x}x^{\prime\prime}<_{x}e$. However this implies (ii), a contradiction. ∎
It is shown in [12] that series-parallel positroids are closed under series-parallel connections, with appropriate ordering on the elements. We extend these results by showing that positroids on non-crossing subsets are closed under series-parallel connections and $2$-sums.
Proposition 2.22.
Let $P$ and $Q$ be positroids on subsets of a densely totally ordered set $X$ such that $\lvert E(P)\rvert\geq 2$, $\lvert E(Q)\rvert\geq 2$, $E(P)\cap E(Q)=\{e\}$, and $E(P)$ and $E(Q)$ are non-crossing subsets of $X$. Then
  • (i)
    $S(P,Q)$ is a positroid.
  • (ii)
    $P(P,Q)$ is a positroid.
  • (iii)
    $P\oplus_{2}Q$ is a positroid.
Proof.
By Observation 2.21, without loss of generality, suppose that $\forall x\in E(P),\forall y\in E(Q)$, $x\leq_{x}e\leq_{x}y$.
By hypothesis $P$ and $Q$ are both positroids, so there exist real matrices $A_{P}$ and $A_{Q}$, with all nonnegative maximal minors, that represent $P$ and $Q$ respectively. As adding a scalar multiple of one row to another row in a square matrix does not change the value of the determinant, we may take $A_{P}$, $A_{Q}$ and $A_{S(P,Q)}$ to be as follows
$A_{P}{=}\left(\begin{array}[]{@{}c|c@{}}A_{P\setminus e}&\begin{matrix}0\\ \vdots\\ 0\\ 1\end{matrix}\end{array}\right),\,A_{Q}{=}\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}1\\ 0\\ \vdots\\ 0\end{matrix}&A_{Q\setminus e}\end{array}\right),\,A_{S(P,Q)}{=}\left(\begin{array}[]{@{}c|c|c@{}}A_{P\setminus e}&\begin{matrix}0\\ \vdots\\ 0\\ 1\end{matrix}&0\\ \hline\cr 0&\begin{matrix}1\\ 0\\ \vdots\\ 0\end{matrix}&A_{Q\setminus e}\end{array}\right).$
The maximal minors of $A_{S(P,Q)}$ are products of maximal minors of $A_{P}$ and $A_{Q}$, which are all nonnegative, so all maximal minors of $A_{S(P,Q)}$ are nonnegative. Therefore, $S(P,Q)$ is a positroid. Positroids are closed under taking duals, so $P(P,Q)=[S(P^{*},Q^{*})]^{*}$ [11, Proposition 7.1.14] is a positroid. As positroids are closed under deletion, $P(P,Q)\setminus e=P\oplus_{2}Q$ is a positroid as well. ∎
Observation 2.23.
Let $P$ and $Q$ be positroids on subsets of a totally ordered set $X$ such that $\lvert E(P)\rvert=2$, $\lvert E(Q)\rvert\geq 2$, and $E(P)\cap E(Q)=\{e\}$. Then $P\oplus_{2}Q$ is a positroid if and only if relabelling $e$ in $Q$ by $E(P)\setminus\{e\}$ is an order preserving matroid isomorphism.
Proof.
Let $E(P)=\{e,p\}$ and $M=P\oplus_{2}Q$. Then $M$ can be obtained from $Q$ by relabelling $e$ with $p$. Thus, $M$ is a positroid if and only if there is an order preserving matroid isomorphism from $M$ to $Q$. ∎
We have established that the $2$-sum of two non-crossing positroids is a positroid, but every $2$-sum decomposition of a positroid need not be into positroids. However, we show that for every $2$-connected positroid that is not $3$-connected, there exists a $2$-sum decomposition into positroids.
Lemma 2.24.
Let $M=P_{1}\oplus_{2}Q_{1}$ be a $2$-connected positroid on a subset of a densely totally ordered set $X$ such that $\lvert E(P_{1})\rvert\geq 3$, $\lvert E(Q_{1})\rvert\geq 3$, and $E(P_{1})\cap E(Q_{1})=\{e_{1}\}$. Then there exist positroids $P_{2}$ obtained from $P_{1}$ and $Q_{2}$ obtained from $Q_{1}$ by relabelling $e_{1}$ with $e_{2}$ in $P_{1}$ and $Q_{1}$ respectively, such that $M=P_{2}\oplus_{2}Q_{2}$.
Proof.
Let $x\in E(P_{1})\setminus\{e_{1}\}$ be fixed. Let $z=\max_{\leq_{x}}\{p\in E(P_{1})\setminus\{e_{1}\}:\forall y\in E(Q_{1})\setminus\{e_{1}\},p\leq_{x}y\}$. Take $e_{2}\in X\setminus\{z\}$ such that for all $y\in E(Q_{1})\setminus\{e_{1}\}$, $z<_{z}e_{2}<_{z}y$. Then, $P_{2}$ and $Q_{2}$ are obtained from $P_{1}$ and $Q_{1}$ by relabelling $e_{1}$ with $e_{2}$, and $M=P_{2}\oplus_{2}Q_{2}$. Let $y^{\prime}=\min_{\leq z}\{y\in E(Q_{1})\setminus\{e_{1}\}:z\leq_{z}y\}$. Then, by Proposition 2.3, $M$ contains a minor $P^{\prime}$ such that $P^{\prime}$ can be obtained from $P_{1}$ by relabelling $e_{1}$ with $y^{\prime}$. Furthermore, as positroids are closed under taking minors, $P^{\prime}$ is a positroid. We can obtain $P_{2}$ from $P^{\prime}$ by relabelling $y^{\prime}$ with $e_{2}$, and this gives an order preserving isomorphism up to rotation, so by Proposition 2.20, $P_{2}$ is a positroid. By symmetry, we have that $Q_{2}$ is also a positroid. ∎
For any $2$-connected positroid $P$, we can take its canonical tree decomposition and appropriately label the edges of the tree $T$ so that any edge-cut of $T$ gives a $2$-sum decomposition of $P$ into non-crossing positroids. To prove this, we first need the following connectivity result.
Proposition 2.25 (Proposition 8.2.8 in [11]).
If $e$ is an element of an $n$-connected matroid $M$, then, provided $\lvert E(M)\rvert\geq 2(n-1)$, both $M\setminus e$ and $M/e$ are $(n-1)$-connected.
Proposition 2.26.
Let $M$ be a $2$-connected positroid on a subset of a densely totally ordered set $X$. Then $M$ has a canonical tree decomposition $T$ such that each $M^{\prime}\in V(T)$ is a positroid. Furthermore, for any edge $e$ in $T$, let $T_{P}$ and $T_{Q}$ be the connected components of $T\setminus\{e\}$. Then the corresponding matroids $P$ and $Q$ are two positroids on non-crossing ground-sets.
Proof.
Let $M$ be a $2$-connected positroid on a subset of a densely totally ordered set $X$. By Theorem 2.6, there exists a canonical tree decomposition $T$ of $M$. Every edge $e_{i}$ of $T$ represents $M$ as a 2-sum $P_{i}\oplus_{2}Q_{i}$. We label each edge $e_{i}$ according to Lemma 2.24, such that every edge splits $T$ into two positroids.
We claim that in addition, every edge of $T$ now cuts $T$ into two non-crossing ground-sets. Suppose by way of contradiction that there exists an edge $e$ of $T$ that cuts $T$ into two positroids $P$ and $Q$, such that $E(P)$ and $E(Q)$ are crossing. Then there exist distinct $a,b,c,d$ such that $a,c\in E(P)$, $b,d\in E(Q)$ and $a<_{a}b<_{a}c<_{a}d$. Suppose that $E(P)\cap E(Q)=e\in\{a,b,c,d\}$. Then by construction, as shown in the proof of Lemma 2.24, there exists some $z\in E(P)-e$ and $y^{\prime}\in E(Q)\setminus\{e\}$, such that $z,e,y$ are consecutive in the cyclic ordering of $E(P)\cup E(Q)$. Therefore, |$|\{a,b,c,d\}\cap\{z,e,y^{\prime}\}|\leq 1$, and $z$, $e$ or $y^{\prime}$ are interchangeable in any crossing pattern. So, we may restrict ourselves to the case where $e\notin\{a,b,c,d\}$.
Let $M_{P}$ and $M_{Q}$ be the respective vertices of the edge $e$ in $T$. By the properties of canonical tree decompositions (Definition 2.5), it is not possible for $M_{P}$ and $M_{Q}$ to both be circuits or cocircuits. We consider two cases: one of $M_{P}$ or $M_{Q}$ is 3-connected, or one is a circuit and one is a cocircuit. Note that a 2-connected matroid has the property that any two of its elements lie on a common circuit and on a common cocircuit.
  • (a)
    Suppose, without loss of generality, that $M_{P}$ is 3-connected and $|E(M_{P})|>3$. Then $M_{P}$ is neither a circuit nor a cocircuit. Then by Proposition 2.25, $M_{P}\setminus e$ and $M_{P}/e$ are both $2$-connected, and therefore $P\setminus e$ and $P/e$ are both $2$-connected. It is known that if a matroid $N$ is 2-connected and $p\in E(N)$, then $N\setminus p$ or $N/p$ is $2$-connected. Therefore, either $Q\setminus e$ or $Q/e$ is $2$-connected. If $Q\setminus e$ is 2-connected, then there is a circuit $C$ in $Q\setminus e$ such that $b,d\in C$, and a cocircuit $C^{*}$ in $P\setminus e$ such that $a,c\in C^{*}$. Since $C$ and $C^{*}$ are disjoint, and both lie in $M$, this is a contradiction. If $Q/e$ is $2$-connected, we similarly find a cocircuit $C^{*}$ in $Q/e$ and a circuit $C$ in $P/e$, both of which are in $M$.
  • (b)
    Suppose, without loss of generality, that $M_{P}$ is a circuit and $M_{Q}$ a cocircuit. Then, $(P/e)^{*}=P^{*}\setminus e$ and $Q\setminus e$ are 2-connected. Therefore, there is a cocircuit $C^{*}$ such that
    $a,b\in C^{*}\in\mathcal{C}^{*}(P/e)\subseteq\mathcal{C}^{*}(M),$
    and there is a circuit $C$, disjoint from $C^{*}$, such that
    $c,d\in C\in\mathcal{C}(Q\setminus e)\subseteq\mathcal{C}(M).$
    Since $C$ and $C^{*}$ are disjoint, and both lie in $M$, this is a contradiction.
Taking Propositions 2.26 and 2.22 together we obtain the following characterization of $2$-sums of positroids.
Theorem 2.27.
Let $M$ be a $2$-connected matroid on a subset of a densely totally ordered set $X$. Then $M$ is a positroid if and only if there exists a canonical tree decomposition $T$ with the following properties. Every vertex of $T$ is a positroid, and cutting $T$ along any edge $e$ results in two 2-connected positroids on non-crossing ground-sets.
2.3. Related combinatorial objects
In [14], Postnikov presented a number of classes of combinatorial objects that are in bijection with positroids. We restrict our consideration to decorated permutations, Grassmann necklaces, and move-equivalent classes of reduced plabic graphs, as these are useful for studying graphic positroids.
2.3.1. Decorated permutations
A decorated permutation on a totally ordered set $X$ is a pair $\pi^{:}=(\pi,col)$ consisting of a permutation $\pi:X\to X$ and a function
$col:\{\mbox{fixed points of }\pi\}\to\{-1,1\}.$
Informally, a decorated permutation is a bijection on a totally ordered set that admits two types of fixed points. For a fixed decorated permutation $\pi^{:}=(\pi,col)$, we define $(\pi^{:})^{-1}=(\pi^{-1},-col)$. For a fixed point $i_{k}$ of $\pi^{:}$, we denote
$\pi^{:}(i_{k})=\begin{cases}\overline{i_{k}},&\text{if }col(i_{k})=-1\\ \underline{i_{k}},&\text{if }col(i_{k})=1.\end{cases}$
Let $i_{j},i_{k}\in\{i_{1}<i_{2}<\cdots<i_{n}\}=X$ and $Y\subseteq X$. We take $\min_{i_{k}}(Y)$ to be the minimum element of $Y$ with respect to the the total order $\leq_{i_{k}}$. We take the cyclic interval $[i_{k},i_{j}]$ to be the set of all elements $i_{t}\in X$ such that $i_{k}\leq_{i_{k}}i_{t}\leq_{i_{k}}i_{j}$.
[…]
$\pi^{:}_{M}(i_{k}):=\begin{cases}\overline{i_{k}},&\text{if }i_{k}\text{ is a coloop}\\ \underline{i_{k}},&\text{if }i_{k}\text{ is a loop}\\ \min_{\leq i_{k}}\left\{i_{j}\in X:i_{k}\in\operatorname{cl}\big{(}[i_{k+1},i_{j}]\big{)}\right\},&\text{otherwise.}\end{cases}$
[…]
Corollary 2.28 (Corollary 13 in [9]).
For any ordered matroid $M$, we have that $\pi^{:}_{M}=(\pi^{:}_{M^{*}})^{-1}$.
We now consider how this map interacts with dihedral action on $X=\{i_{1}<i_{2}<\cdots<i_{n}\}$. Let $\pi^{:}$ be a decorated permutation on $X$ and let $D_{X}$ be the group of bijections on $X$ under composition, generated by $r(i_{k})=i_{k+1}$ and $s(i_{k})=i_{n-k+1}$ where the indices are taken modulo $n$. We define the group action of $D_{X}$ on the set {decorated permutations on $X$} as follows, for all $\omega\in D_{X}$
$(\omega\cdot\pi^{:})(x):=\begin{cases}\overline{\omega(x)},&\text{if }\pi^{:}(x)=\overline{x}\\ \underline{\omega(x)},&\text{if }\pi^{:}(x)=\underline{x}\\ \omega(\pi^{:}(x)),&\text{otherwise.}\end{cases}$
[…]
Remark 2.29.
For any matroid $M$ on $\{i_{1}<i_{2}<\cdots<i_{n}\}$, we have that $r\cdot\pi^{:}_{M}=\pi^{:}_{r\cdot M}$.
[…]
Lemma 2.30.
For any matroid $M$ on $\{i_{1}<i_{2}<\cdots<i_{n}\}$, we have that
$\pi^{:}_{s\cdot M}(s(i_{k}))=\begin{cases}\overline{s(i_{k})},&\text{if }\pi^{:}_{M}(i_{k})=\overline{s(i_{k})}\\ \underline{s(i_{k})},&\text{if }\pi^{:}_{M}(i_{k})=\underline{s(i_{k})}\\ s((\pi^{:}_{M}(i_{k}))^{-1}),&\text{otherwise.}\end{cases}$
Proof.
Let $i_{k}\in X$. Suppose that $i_{k}$ is a coloop in $M$. Then $s(i_{k})$ is a coloop in $s\cdot M$ and $\pi^{:}_{s\cdot M}(s(i_{k}))=\overline{s(i_{k})}$. Suppose instead that $i_{k}$ is a loop in $M$. Then $s(i_{k})$ is a loop in $s\cdot M$ and $\pi^{:}_{s\cdot M}(s(i_{k}))=\underline{s(i_{k})}$.
Now suppose that $i_{k}$ is neither a loop nor a coloop in $M$. Then $s(i_{k})$ is neither a loop nor a coloop in $s\cdot M$. Furthermore,
$X=\{s(i_{k})<_{s(i_{k})}\cdots<_{s(i_{k})}s(i_{1})<_{s(i_{k})}s(i_{n})<_{s(i_{k})}\cdots<_{s(i_{k})}s(i_{k+1})\},$
so
$\pi^{:}_{s\cdot M}(s(i_{k}))$ $=\min_{\leq_{s(i_{k})}}\left\{s(i_{j})\in X:s(i_{k})\in\operatorname{cl}([s(i_{k-1}),s(i_{j})]\right\}$
$=s((\pi^{:}_{M}(i_{k}))^{-1}).$
When an ordered matroid $M$ is loopless and coloopless, we obtain a nicer characterization of $\pi^{:}_{s\cdot M}$.
Corollary 2.31.
For any loopless and coloopless matroid $M$, we have that $\pi^{:}_{s\cdot M}=s\cdot(\pi^{:}_{M})^{-1}$.
Let us now consider the interactions between decorated permutations and the matroid operations of direct sums, series-parallel connections, and $2$-sums. Constructions similar to the ones we present have been given in [12, 7]. Ours differ by considering decorated permutations on any finite subset of a totally ordered set, not necessarily $[n]$ for some integer $n$. This is intended to deal with the technical challenges of decompositions of decorated permutations induced by direct sum and $2$-sum decompositions of their corresponding matroids. We begin by defining the following binary operation on decorated permutations.
Definition 2.32.
Let $X$ be a totally ordered set and let $\pi^{:}_{1},\pi^{:}_{2}$ be decorated permutations on $X_{1}$ and $X_{2}$ respectively, such that $X_{1}$ and $X_{2}$ are disjoint subsets of $X$. We define the disjoint union of $\pi^{:}_{1}$ and $\pi^{:}_{2}$, which we denote by $\pi^{:}_{1}\sqcup\pi^{:}_{2}$, as the decorated permutation given by
$(\pi^{:}_{1}\sqcup\pi^{:}_{2})(k)=\begin{cases}\pi^{:}_{1}(k),&\text{if }k\in X_{1}\\ \pi^{:}_{2}(k),&\text{if }k\in X_{2}.\end{cases}$
Remark 2.33.
Let $M$ and $N$ be matroids on disjoint subsets of a totally ordered set $X$. Then,
$\pi^{:}_{M\oplus N}=\pi^{:}_{M}\sqcup\pi^{:}_{N}.$
This characterization of the decorated permutations of direct sums of matroids using disjoint unions is analogous to the work of Moerman and Williams in [7, Section 4] and Parisi, Shermann-Bennett, and Williams in [12, Section 12] using direct sums of permutations. We now turn our attention to the decorated permutations of the $2$-sums of matroids.
Lemma 2.34.
Let $M$ and $N$ be matroids on subsets of $X=\{i_{1}<i_{2}<\cdots<i_{n}\}$ such that $\lvert E(M)\rvert\geq 2$, $\lvert E(N)\rvert\geq 2$, and $E(M)\cap E(N)=\{i_{e}\}$. Suppose that $E(M)$ and $E(N)$ are non-crossing subsets of $X$. Suppose further that $i_{e}$ is a loop or coloop of $M$ or $N$, then $\pi^{:}_{M\oplus_{2}N}$ is given by
$\pi^{:}_{M\oplus_{2}N}=\pi^{:}_{M\setminus i_{e}\oplus N\setminus i_{e}}.$
Suppose instead that $i_{e}$ is neither a loop nor a coloop of $M$ or $N$, then $\pi^{:}_{M\oplus_{2}N}$ is given by
$\pi^{:}_{M\oplus_{2}N}(i_{k}):=\begin{cases}\pi^{:}_{M}(i_{k}),&\text{if }i_{k}\in E(M)-E(N),\pi^{:}_{M}(i_{k})\neq i_{e}\\ \pi^{:}_{N}(i_{e}),&\text{if }i_{k}\in E(M)-E(N),\pi^{:}_{M}(i_{k})=i_{e}\\ \pi^{:}_{N}(i_{k}),&\text{if }i_{k}\in E(N)-E(M),\pi^{:}_{N}(i_{k})\neq i_{e}\\ \pi^{:}_{M}(i_{e}),&\text{if }i_{k}\in E(N)-E(M),\pi^{:}_{N}(i_{k})=i_{e}.\end{cases}$
Proof.
By Observation 2.21, without loss of generality, suppose that $\forall i_{x}\in E(M)$, $\forall i_{y}\in E(N)$, $i_{x}\leq_{i_{x}}i_{e}\leq_{i_{x}}i_{y}$.
Suppose that $i_{e}$ is a loop or coloop in $M$ or $N$, then $M\oplus_{2}N=M\setminus i_{e}\oplus N\setminus i_{e}$. Thus $\pi^{:}_{M\oplus_{2}N}=\pi^{:}_{M\setminus i_{e}\oplus N\setminus i_{e}}$.
Now suppose that $i_{e}$ is neither a loop nor a coloop of $M$ or $N$. Suppose further that $i_{k}\in E(M)\setminus E(N)$ and $\pi^{:}_{M}(i_{k})\neq i_{e}$. Then
$\pi^{:}_{M\oplus_{2}N}(i_{k})$ $=\min_{\leq_{i_{k}}}\{i_{j}\in E(M\oplus_{2}N):i_{k}\in\operatorname{cl}_{M\oplus_{2}N}\big{(}[i_{k+1},i_{j}]\big{)}\}$
$=\min_{\leq_{i_{k}}}\{i_{j}\in E(M\setminus i_{e}):i_{k}\in\operatorname{cl}_{M}\big{(}[i_{k+1},i_{j}]\big{)}\}$
$=\pi^{:}_{M}(i_{k}).$
Suppose instead that $i_{k}\in E(M)\setminus E(N)$ and $\pi^{:}_{M}(i_{k})=i_{e}$. Then
$\pi^{:}_{M\oplus_{2}N}(i_{k})$ $=\min_{\leq_{i_{k}}}\{i_{j}\in E(M\oplus_{2}N):i_{k}\in\operatorname{cl}_{M\oplus_{2}N}\big{(}[i_{k+1},i_{j}]\big{)}\}$
$=\min_{\leq_{i_{k}}}\{i_{j}\in E(N\setminus i_{e}):i_{k}\in\operatorname{cl}_{N}\big{(}[i_{e},i_{j}]\big{)}\}$
$=\pi^{:}_{N}(i_{e}).$
The remaining cases follow a similar argument. ∎
The decorated permutations of $2$-sums can analogously be defined in the language of amalgamations of decorated permutations, as shown by Moerman and Williams in [7, Section 4]. We complete this section by considering the decorated permutations of series-parallel connections. Parisi, Shermann-Bennett, and Williams have presented similar constructions in [12, Section 12].
Lemma 2.35.
Let $M$ and $N$ be $2$-connected matroids on subsets of $X=\{i_{1}<i_{2}<\cdots<i_{n}\}$ such that $\lvert E(M)\rvert\geq 2$, $\lvert E(N)\rvert\geq 2$, and $E(M)\cap E(N)=\{i_{e}\}$. Suppose that $E(M)$ and $E(N)$ are non-crossing subsets of $X$. Then $\pi^{:}_{P(M_{1},M_{2})}$ is given by
$\pi^{:}_{P(M,N)}(i_{k}):=\begin{cases}\pi^{:}_{M}(i_{k}),&\text{if }i_{k}\in E(M)\setminus E(N)\\ \pi^{:}_{N}(i_{e}),&\text{if }i_{k}=i_{e}\\ \pi^{:}_{N}(i_{k}),&\text{if }i_{k}\in E(N)\setminus E(M),\pi^{:}_{N}(i_{k})\neq i_{e}\\ \pi^{:}_{M}(i_{e}),&\text{if }i_{k}\in E(N)\setminus E(M),\pi^{:}_{N}(i_{k})=i_{e},\end{cases}$
and $\pi^{:}_{S(M,N)}$ is given by
$\pi^{:}_{S(M,N)}(i_{k}):=\begin{cases}\pi^{:}_{M}(i_{k}),&\text{if }i_{k}\in E(M),\pi^{:}_{M}(i_{k})\neq i_{e}\\ \pi^{:}_{N}(i_{e}),&\text{if }i_{k}\in E(M),\pi^{:}_{M}(i_{k})=i_{e}\\ \pi^{:}_{N}(i_{k}),&\text{if }i_{k}\in E(N)\setminus E(M),\pi^{:}_{N}(i_{k})\neq i_{e}\\ \pi^{:}_{M}(i_{e}),&\text{if }i_{k}\in E(N)\setminus E(M),\pi^{:}_{N}(i_{k})=i_{e}.\end{cases}$
Proof.
By Observation 2.21, there is an element $i_{r}$ such that $E(M)=[i_{r},\dots,i_{e}]$ and $E(N)=[i_{e},\dots,i_{r-1}]$. We assume $r=1$ without loss of generality.
For $1\leq k<e$, we have
$\pi^{:}_{P(M,N)}(i_{k})$ $=\min_{\leq_{i_{k}}}\left\{i_{j}\in E(P(M,N)):i_{k}\in\operatorname{cl}_{P(M,N)}\big{(}[i_{k+1},i_{j}]\big{)}\right\}$
$=\min_{\leq_{i_{k}}}\left\{i_{j}\in E(M):i_{k}\in\operatorname{cl}_{M}\big{(}[i_{k+1},i_{j}]\big{)}\right\}$
$=\pi^{:}_{M}(i_{k}).$
In the remainder of this proof, we omit the first equality that is simply the definition of the decorated permutation. We have,
$\pi^{:}_{P(M,N)}(i_{e})=\min_{\leq_{i_{e}}}\left\{i_{j}\in E(N):i_{e}\in\operatorname{cl}_{N}\big{(}[i_{e+1},i_{j}]\big{)}\right\}=\pi^{:}_{N}(i_{e}).$
Now suppose that $1\leq k\leq e$ and $\pi^{:}_{M}(i_{k})\neq i_{e}$. Then
$\pi^{:}_{S(M,N)}(i_{k})=\min_{\leq_{i_{k}}}\left\{i_{j}\in E(M):i_{k}\in\operatorname{cl}_{M}\big{(}[i_{k+1},i_{j}]\big{)}\right\}=\pi^{:}_{M}(i_{k}).$
Suppose instead that $1\leq k\leq e$ and $\pi^{:}_{M}(i_{k})=i_{e}$. Then
$\pi^{:}_{S(M,N)}(i_{k})=\min_{\leq_{i_{e}}}\left\{i_{j}\in E(N):i_{e}\in\operatorname{cl}_{N}\big{(}[i_{e+1},i_{j}]\big{)}\right\}=\pi^{:}_{N}(i_{e}).$
Now suppose that $e<k\leq n$ and $\pi^{:}_{N}(i_{k})\neq i_{e}$. Then
$\pi^{:}_{P(M,N)}(i_{k})$ $=\min_{\leq_{i_{k}}}\left\{i_{j}\in E(N):i_{k}\in\operatorname{cl}_{N}\big{(}[i_{k+1},i_{j}]\big{)}\right\}=\pi^{:}_{N}(i_{k}),$
$\pi^{:}_{S(M,N)}(i_{k})$ $=\min_{\leq_{i_{k}}}\left\{i_{j}\in E(N):i_{k}\in\operatorname{cl}_{N}\big{(}[i_{k+1},i_{j}]\big{)}\right\}=\pi^{:}_{N}(i_{k}).$
Suppose instead that $e<k\leq n$ and $\pi^{:}_{N}(i_{k})=i_{e}$. Then
$\pi^{:}_{P(M,N)}(i_{k})$ $=\min_{\leq_{i_{k}}}\left\{i_{j}\in E(M):i_{e}\in\operatorname{cl}_{M}\big{(}[i_{r},i_{j}]\big{)}\right\}=\pi^{:}_{M}(i_{e}),$
$\pi^{:}_{S(M,N)}(i_{k})$ $=\min_{\leq_{i_{e}}}\left\{i_{j}\in E(M):i_{e}\in\operatorname{cl}_{M}\big{(}[i_{e+1},i_{j}]\big{)}\right\}=\pi^{:}_{M}(i_{e}).$
2.3.2. Grassmann necklaces
Definition 2.36.
A Grassmann necklace on a totally ordered set $X=\{i_{1}<i_{2}<\cdots<i_{n}\}$ is a sequence $\mathcal{J}=(J_{i_{1}},\ldots,J_{i_{n}})$ of subsets $J_{i_{r}}\subseteq X$ given by
$J_{i_{k+1}}=\begin{cases}(J_{i_{k}}\setminus\{i_{k}\})\cup i_{j},&\text{if }i_{k}\in J_{i_{k}}\\ J_{i_{k+1}}=J_{i_{k}},&\text{otherwise,}\end{cases}$
where the indices are taken modulo $n$.
[…]
$\pi^{:}_{\mathcal{J}}(i_{k})=\begin{cases}i_{j},&\text{if }J_{i_{k+1}}=(J_{i_{k}}\setminus\{i_{k}\})\cup i_{j},i_{j}\neq i_{k}\\ \overline{i_{k}},&\text{if }J_{i_{k+1}}=J_{i_{k}},i_{k}\in J_{i_{k}}\\ \underline{i_{k}},&\text{if }J_{i_{k+1}}=J_{i_{k}},i_{k}\notin J_{i_{k}}.\end{cases}$
Let $\pi^{:}$ be a decorated permutation on $X$ and $i_{k}\in X$, then the Grassmann necklace associated to $\pi^{:}$ is $\mathcal{J}(\pi^{:})=(J_{i_{1}},\ldots,J_{i_{n}})$ and is given by
$J_{i_{k}}=\left\{i_{j}:1\leq j\leq n,(\pi^{:}(i_{j}))^{-1}>_{i_{k}}i_{j}\mbox{or }\pi^{:}(i_{j})=\overline{i_{j}}\right\}.$
{ordered matroids on $X$}{decorated permutations on $X$}{Grassmann necklaces on $X$}
Figure 1. For a totally ordered set $X$, this diagram commutes.
[…]
2.4. Positroid envelopes, envelope classes, and varieties
In this subsection we provide relevant background on positroid envelopes and positroid varieties. We define positroid envelope classes along with binary operations on the positroid envelope classes that are the analogues of direct sums and $2$-sums of positroids. We then show that the positroid envelope classes have decompositions under these direct sum and $2$-sum operations. The reader is directed to [5], wherein Knutson, Lam, and Speyer originally define positroid envelopes, for more information on positroid envelopes and positroid varieties.
2.4.1. Positroid envelopes
[…]
Theorem 2.37 (Theorem 8 in [10]).
Let $P$ be a matroid on $X=\{i_{1}<i_{2}<\cdots<i_{n}\}$ of rank $k$ and $\mathcal{J}(P)=(J_{i_{1}},J_{i_{2}},\ldots,J_{i_{n}})$ its corresponding Grassmann necklace. Then $P$ is a positroid if and only if
$\mathcal{B}(P)=\bigcap^{n}_{j=1}\left\{B={X\choose k}:J_{i_{j}}\leq_{i_{j}}B\right\}.$
This immediately implies that if $P$ is a positroid and $M$ an ordered matroid such that $\mathcal{J}(M)=\mathcal{J}(P)$, then $\mathcal{B}(M)\subseteq\mathcal{B}(P)$. Furthermore, we say that $P$ is the positroid envelope of $M$ [5]. As any matroid $M$ whose positroid envelope is $P$ has the same rank as $P$, we obtain the following.
Corollary 2.38.
Let $M$ be an ordered matroid with positroid envelope $P$. Then $\mathbb{1}:P\to M$ is rank preserving weak map.
The next two corollaries immediately follow from applying results for rank-preserving weak maps to positroid envelopes.
Corollary 2.39.
Let $M$ be an ordered matroid with positroid envelope $P$. Let $N$ be a minor of $M$. Then there exists a minor $Q$ of $P$ such that $\mathbb{1}:Q\to N$ is a rank-preserving weak map.
Proof.
This follows from Corollary 2.38 and Lemma 2.12. ∎
Corollary 2.40.
Let $M$ be an ordered matroid with positroid envelope $P$. Suppose that $P$ is $U^{k}_{n}$-free, then $M$ is $U^{k}_{n}$-free.
Proof.
This follows from Corollaries 2.38 and 2.13. ∎
For any minor $N$ of an ordered matroid $M$, where $Q$ is the positroid envelope of $N$ and $P$ is the positroid envelope of $M$, there exists a minor $K$ of $P$ such that $\mathbb{1}:K\to Q$ is a rank-preserving weak map.
Lemma 2.41.
Let $M$ be an ordered matroid with positroid envelope $P$. Suppose that $N$ is a minor of $M$, and let $Q$ be the positroid envelope of $N$. Then there exists a minor $K$ of $P$ such that $\mathbb{1}:K\to Q$ is a rank-preserving weak map.
Proof.
Without loss of generality, let $Q$ and $N$ be ordered matroids on $X=\{i_{1}<i_{2}<\cdots<i_{n}\}$. By Theorem 2.12, there exists a minor $K$ of $P$ such that $\mathbb{1}:K\to N$ is a rank-preserving weak map and therefore $\mathcal{B}(N)\subseteq\mathcal{B}(K)$. Let $\mathcal{J}(Q)=(J_{i_{1}},J_{i_{2}},\ldots,J_{i_{n}})=\mathcal{J}(N)$ and let $\mathcal{J}(K)=(J^{\prime}_{i_{1}},J^{\prime}_{i_{2}},\ldots,J^{\prime}_{i_{n}})$. Then for all $k\in[n]$, $J^{\prime}_{i_{k}}\leq_{i_{k}}J_{i_{k}}$. By Theorem 2.37,
$\mathcal{B}(Q)$ $=\bigcap^{n}_{k=1}\left\{B={X\choose r}:J_{i_{k}}\leq_{i_{k}}B\right\}$
$\subseteq\bigcap^{n}_{k=1}\left\{B={X\choose r}:J^{\prime}_{i_{k}}\leq_{i_{k}}B\right\}=\mathcal{B}(K).$
Therefore, $\mathbb{1}:K\to Q$ is a rank-preserving weak map. ∎
For any rank-$k$ matroid $M$ on $n$ elements, there exists a unique matroid $N$ isomorphic to the uniform matroid $U^{k}_{n}$ such that $\mathbb{1}:N\to M$ is a rank-preserving weak map. We use this fact, together with Lemma 2.41, to provide sufficient conditions for a positroid to contain a minor isomorphic to a uniform matroid.
Corollary 2.42.
Let $M$ be an ordered matroid with positroid envelope $P$. Suppose that $M$ has a minor $N$ with positroid envelope $Q\cong U^{k}_{n}$. Then $Q$ is a minor of $P$.
Proof.
Let $X$ be the totally ordered ground-set of $Q$. By Lemma 2.41, $P$ contains a minor $K$ such that $\mathbb{1}:K\to Q$ is a rank-preserving weak map. Thus,
$\mathcal{B}(Q)=\left\{{X\choose k}\right\}\subseteq\mathcal{B}(K)\subseteq\left\{{X\choose k}\right\}.$
Therefore, $K=Q$. ∎
We now consider the equivalence classes of ordered matroids where two ordered matroids are equivalent if they share the same positroid envelope. We formalize this in the following definition.
Definition 2.43 (Positroid envelope class).
For a fixed positroid $P$, the envelope class $\Omega_{P}$ is the set of all ordered matroids whose positroid envelope is $P$.
Note that we may equivalently define the positroid envelope classes as follows.
$\Omega_{P}=\{M:\mathcal{J}(M)=\mathcal{J}(P)\}=\{M:\pi^{:}_{M}=\pi^{:}_{P}\}.$
By Corollary 2.28, we arrive at the following duality result for positroid envelope classes.
Proposition 2.44.
For any positroid envelope class $\Omega_{P}$, we have
$\Omega_{P^{*}}=\{M^{*}:M\in\Omega_{P}\}.$
For a fixed finite, totally ordered set $X$, with $D_{X}$ the group isomorphic to the dihedral group on $2|X|$, we define the group action of $D_{X}$ on {positroid envelope classes on $X$} as follows. For $\omega\in D_{X}$,
$\omega\cdot\Omega_{P}:=\{\omega\cdot M:M\in\Omega_{P}\}.$
Dihedral action on positroid envelope classes has the following nice property.
Proposition 2.45.
For a positroid envelope class $\Omega_{P}$ and $\omega\in D_{X}$, we have $\omega\cdot\Omega_{P}=\Omega_{\omega\cdot P}$.
Proof.
Let $\omega\in D_{X}$, then $\omega\cdot\Omega_{P}=\{\omega\cdot M:\pi^{:}_{M}=\pi^{:}_{P}\}$. It follows from Remark 2.29 and Lemma 2.30 that
$\omega\cdot\Omega_{P}=\{\omega\cdot M:\pi^{:}_{M}=\pi^{:}_{P}\}=\{M:\pi^{:}_{\omega\cdot M}=\pi^{:}_{\omega\cdot P}\}=\Omega_{\omega\cdot P}.$
If the decorated permutation corresponding to a positroid $P$ consists of a single cycle, then every matroid in $\Omega_{P}$ is $2$-connected.
Proposition 2.46.
Let $P$ be a positroid on $X=\{i_{1}<i_{2}<\cdots<i_{n}\}$. Suppose that $\pi^{:}_{P}$ consists of a single cycle. Then $\forall M\in\Omega_{P}$, $M$ is $2$-connected.
Proof.
Let $i_{k}\in X$, then $X=\{i_{k},\pi^{:}(i_{k}),(\pi^{:}(i_{k}))^{2},\ldots,(\pi^{:}(i_{k}))^{n-1}\}$. Therefore, $\forall 0\leq j<n-1$, $(\pi^{:}(i_{k}))^{j}$ and $(\pi^{:}(i_{k}))^{j+1}$ lie on a common circuit. Thus $M$ is $2$-connected. ∎
We can now show that if a decorated permutation contains two cycles on crossing subsets, then the corresponding positroid contains a minor isomorphic to $U^{2}_{4}$ and is therefore non-binary.
Lemma 2.47.
Let $P$ be a positroid and $\{a<b<c<d\}\subseteq E(P)$. Suppose that $\pi^{:}_{P}$ contains two disjoint cycles $\pi^{:}_{1}$ and $\pi^{:}_{2}$, on $X_{1}$ and $X_{2}$ respectively, such that $a,c\in X_{1}$ and $b,d\in X_{2}$. Then $P$ contains a minor isomorphic to $U^{2}_{4}$.
Proof.
Take $\pi^{:}_{P}=\bigsqcup^{n}_{k=1}\pi^{:}_{i}$, where $\{\pi^{:}_{k}\}^{n}_{k=1}$ is the set of disjoint cycles in $\pi^{:}$. Then for all $k\in[n]$, $\pi^{:}_{k}$ is a decorated permutation on $X_{k}\subset E(P)$, where $\{X_{k}\}^{n}_{k=1}$ is a partition of $E(P)$, and there exist positroids $\{P_{k}\}^{n}_{k=1}$ such that $\pi^{:}_{P_{k}}=\pi^{:}_{k}$. Thus, $M=\bigoplus^{n}_{k=1}P_{k}\in\Omega_{P}$. By restricting to $E(P_{1}\oplus P_{2})$, we obtain $P_{1}\oplus P_{2}$ as a minor of $M$, where $E(P_{1})=X_{2}$, $E(P_{2})=X_{2}$, $\pi^{:}_{P_{1}}=\pi^{:}_{1}$, and $\pi^{:}_{P_{2}}=\pi^{:}_{2}$. As $\pi^{:}_{1}$ and $\pi^{:}_{2}$ both consist of single cycles, by Proposition 2.46, both $P_{1}$ and $P_{2}$ are $2$-connected matroids. Then there exist circuits $C_{1}\in\mathcal{C}(P_{1})$ and $C_{2}\in\mathcal{C}(P_{2})$ such that $a,c\in C_{1}$ and $b,d\in C_{2}$. Therefore, $N_{1}=(P_{1}\rvert C_{1})/(C_{1}\setminus\{a,c\})\cong U^{1}_{2}$ and $N_{2}=(P_{2}\rvert C_{2})/(C_{2}\setminus\{b,d\})\cong U^{1}_{2}$ are minors of $P_{1}$ and $P_{2}$ respectively, so $N_{1}\oplus N_{2}$ is a minor of $P_{1}\oplus P_{2}$. Furthermore, $\pi^{:}_{N_{1}}=(a,c)$ and $\pi^{:}_{N_{2}}=(b,d)$, so by Remark 2.33, $\pi^{:}_{N_{1}\oplus N_{2}}=(a,c)(b,d)$. Hence, there exists a unique positroid $Q$ such that $\pi^{:}_{Q}=(a,c)(b,d)$ and $Q\cong U^{2}_{4}$. Then, by Lemma 2.41, there exists a positroid $K$ that is a minor of $P$ such that $E(K)=E=E(Q)$ and $\mathbb{1}:K\to Q$. Thus,
{($\bigg{\{}{E\choose 2}\bigg{\}}=\mathcal{B}(Q)\subseteq\mathcal{B}(K)\subseteq\bigg{\{}{E\choose 2}\bigg{\}}.$
Therefore, $\mathcal{B}(K)\cong U^{2}_{4}$. ∎
We now present some connectivity results for rank-preserving weak maps that we apply to direct sum decompositions of positroid envelope classes in Proposition 2.52.
Lemma 2.48.
Let $M$ and $N$ be matroids such that $\mathbb{1}:M\to N$ is a rank-preserving weak map. Suppose that for $X\subseteq E$, $(X,E\setminus X)$ is a $k$-separation of $M$. Then $(X,E)$ is a $k$-separation of $N$.
Proof.
As $\mathbb{1}:M\to N$ is a weak map, $\mathcal{I}(N)\subseteq\mathcal{I}(M)$, hence $\forall X\in E$, $r_{N}(X)\leq r_{M}(X)$. By hypothesis, $(X,E)$ is a $k$-separation of $M$, so $\min\{\lvert X\rvert,\lvert E\setminus X\rvert\}\geq k$ and
$k>\lambda_{M}(X_{1})$ $=r_{M}(X_{1})+r_{M}(X_{2})-r_{M}(M)$
$\geq r_{N}(X_{1})+r_{N}(X_{2})-r_{N}(N)=\lambda_{N}(X_{1}).$
Therefore, $(X,E\setminus X)$ is a $k$-separation of $N$. ∎
It follows that if the image of a rank-preserving weak map is $k$-connected, then the pre-image is $k$-connected. Furthermore, if an ordered matroid is $k$-connected, then its positroid envelope is $k$-connected.
Corollary 2.49.
Let $M$ and $N$ be matroids such that $\mathbb{1}:M\to N$ is a rank-preserving weak map. If $N$ is $k$-connected, then $M$ is $k$-connected.
Corollary 2.50.
Let $M$ be an ordered matroid with positroid envelope $P$. If $M$ is $k$-connected, then $P$ is $k$-connected.
We define the direct sum of positroid envelope classes, a binary operation on positroid envelope classes that is the analogue of direct sums of matroids. We then use this construction to present a direct sum decomposition of positroid envelope classes.
Definition 2.51.
Let $M$ and $N$ be matroids on subsets of a totally ordered set $X$, with positroid envelopes $P$ and $Q$ respectively. Suppose that $E(M)$ and $E(N)$ are disjoint. Then we define the direct sum of envelope classes $\Omega_{P}$ and $\Omega_{Q}$ as
$\Omega_{P}\oplus\Omega_{Q}=\{M^{\prime}\oplus N^{\prime}:M^{\prime}\in\Omega_{P},N^{\prime}\in\Omega_{Q}\}.$
Proposition 2.52.
Let $P$ and $Q$ be positroids on disjoint, non-crossing subsets $X_{1}$ and $X_{2}$ of a totally ordered set $X$. Then
$\Omega_{P\oplus Q}=\Omega_{P}\oplus\Omega_{Q}.$
Proof.
Suppose that $M_{1}\in\Omega_{P}$ and $M_{2}\in\Omega_{Q}$. Then, by Remark 2.33,
$\pi^{:}_{M_{1}\oplus M_{2}}=\pi^{:}_{M_{1}}\sqcup\pi^{:}_{M_{2}}=\pi^{:}_{P\oplus Q}.$
As $M_{1}\in\Omega_{P}$ and $M_{2}\in\Omega_{Q}$ are arbitrary, $\Omega_{P}\oplus\Omega_{Q}\subseteq\Omega_{P\oplus Q}$.
Suppose instead that $M\in\Omega_{P\oplus Q}$. Then, by Lemma 2.48, we have that $(E(P),E(Q))$ is a $1$-separation of $M$. Let $M_{1}=M|E(P)$ and $M_{2}=M|E(Q)$, then $M=M_{1}\oplus M_{2}$. Furthermore, $\pi^{:}_{M_{1}}=\pi^{:}_{P}$ and $\pi^{:}_{M_{2}}=\pi^{:}_{Q}$. As $M\in\Omega_{P\oplus Q}$ is arbitrary, $\Omega_{P\oplus Q}\subseteq\Omega_{P}\oplus\Omega_{Q}$. Therefore, $\Omega_{P\oplus Q}=\Omega_{P}\oplus\Omega_{Q}$. ∎
2.4.2. Positroid varieties
For a field $\mathbb{F}$ and nonnegative integers $k<n$, the Grassmannian $\operatorname{Gr}(k,\mathbb{F}^{n})$ is the collection of all $k$-dimensional linear subspaces of the vector space $\mathbb{F}^{n}$. Let $V\in\operatorname{Gr}(k,\mathbb{F}^{n})$, then $V$ is the row space of a full-rank $k\times n$ matrix $A$. For two distinct full-rank $k\times n$ matrices $A$ and $B$ with entries over $\mathbb{F}$, whose row spaces are both $V$, we have $M(A)=M(B)$. Thus, we may unambiguously define the ordered matroid $M_{V}=M(A)$ where $A$ is any matrix whose row space is the linear subspace $V$. For a rank-$k$ ordered matroid $M$ on $n$ elements, we define the matroid stratum in $\operatorname{Gr}(k,\mathbb{F}^{n})$ as
$S_{M}(\mathbb{F}):=\{V\in\operatorname{Gr}(k,\mathbb{F}^{n}):M_{V}=M\}.$
The totally nonnegative Grassmannian $\operatorname{Gr}^{\geq 0}(k,\mathbb{R}^{n})$ consists of the points in the Grassmannian $\operatorname{Gr}(k,\mathbb{R}^{n})$ that can be realized by matrices with entries over $\mathbb{R}$ with all nonnegative maximal minors.The positroid cell corresponding to the ordered matroid $M$ is defined as
$S_{M}^{\geq 0}(\mathbb{R}):=S_{M}(\mathbb{R})\cap\operatorname{Gr}^{\geq 0}(k,\mathbb{R}^{n}).$
The positroid stratification of the totally nonnegative Grasssmannian $\operatorname{Gr}^{\geq 0}(k,\mathbb{R}^{n})$ is given by the intersection of the matroid strata with $\operatorname{Gr}(k,\mathbb{R}^{n})$, which is exactly the collection of positroid cells. Note that an ordered matroid is a positroid if and only if its corresponding positroid cell is non-empty.
For a positroid $P$, its open positroid variety is defined as
$\Pi^{\circ}_{\mathbb{F}}(P):=\{V\in\operatorname{Gr}(k,\mathbb{F}^{n}):M_{V}\in\Omega_{P}\}=\bigsqcup_{M\in\Omega_{P}}S_{M},$
and its positroid variety $\Pi_{\mathbb{F}}(P)=\overline{\Pi^{\circ}_{\mathbb{F}}(P)}$.
Corollary 2.53 (Corollary 5.12 in [5]).
Let $P$ be a positroid. Then, as sets,
$\Pi_{\mathbb{R}}(P)=\overline{S_{P}}=\{V\in\operatorname{Gr}(k,\mathbb{R}^{n}):I\notin\mathcal{I}(P)\Rightarrow\Delta_{I}(V)=0\}.$
3. Graph constructions
In this section, we show that for any decorated permutation $\pi^{:}$, there exists a planar graph $G$ such that $\pi^{:}=\pi^{:}_{M(G)}$. Since fixed points of $\pi^{:}$ are represented by self-loops or leaf-edges anywhere in $G$, and distinct cycles of the permutation may be represented by distinct components of $G$, the non-trivial part of this proof is in the following lemma, which deals with representing single $m$-cycles.
Lemma 3.1.
Let $\pi$ be an $m$-cycle on $[m]$. Then, there exists a 2-connected planar graph $G$ such that $\pi=\pi^{:}_{M(G)}$.
Proof.
We create our graph $G$ as follows. Let
$V(G)=\{v_{i}:1\leq i\leq m,\;i<\pi^{-1}(i)\}.$
In the following construction, we consider the element $m$ to be both greater and less than all other elements. For each $1\leq i\leq m$, we let $f_{R}(i)$ be the first element in the list $\pi(i),\pi^{2}(i),\dots,m$ that is less than $i$, and $f_{L}(i)$ the first element in the list $\pi^{-1}(i),\pi^{-2}(i),\dots,m$ that is greater than $i$. Each (directed) edge $i$ goes from $i_{out}=v_{f_{R}(i)}$ to $i_{in}=v_{\pi(f_{L}(i))}$. By construction, these vertices are indeed in the set $V(G)$.
We embed $G$ in the plane as follows. In an $(m+1)\times(m+1)$ grid, put the label $m$ in boxes $(0,m)$ and $(m,0)$. Then, for $11\leq i\leq m-$ 1, put the label $\pi^{i}(m)$ in the box $(i,\pi^{i}(m))$. For $11\leq i\leq m-$ 1, draw each vertex $v_{i}$ stretching from the bottom right corner of $\pi^{-1}(i)$ to the top left corner of $i$. Draw each edge $i$ horizontally from $i_{out}$ to the bottom right corner of the box $i$, then to its top left corner, and horizontally to vertex $i_{in}$. The edge $m$ goes in to the bottom right corner of the box $(0,m)$, loops around outside of the grid, and comes out of the top left corner of $(m,0)$, as shown in Figure 2. By construction, this is a planar embedding of $G$.
We now embed $G^{*}$ by putting the label $m$ in the boxes $(0,0)$ and $(m,m)$, and perform the left-right mirror image of the embedding algorithm. The graph $G^{*}$ has edge set $i^{*}$ for $1\leq i\leq m$ and vertex set
$V(G^{*})=\{v_{i}^{*}:1\leq i\leq m,\;i<\pi(i)\}.$
We first claim that edges of $G$ and $G^{*}$ only cross when an edge $i$ crosses $i^{*}$ in the box that is labelled $i$, for $11\leq i\leq m-$ 1, and the edge $m$ crosses $m^{*}$ outside of the grid. It is easy to see that $m$ and $m^{*}$ have no other crossings. Note that outside of labeled boxes, edges are horizontal on gridlines. Consider the gridline between two rows $k$ and $k+1$, for some $1\leq k<m$. Without loss of generality, suppose that the box $k$ is left of $k+1$. Then all edges of $G$ on the gridline are either left of $k$ or right of $k+1$, while edges of $G^{*}$ on the gridline are right of $k$ and left of $k+1$. Therefore, there are no crossings of edges in $G$ and $G^{*}$ other than those inside labeled boxes.
We argue that $G^{*}$ is indeed the dual of $G$, by showing a bijection between vertices of $G^{*}$ and faces of $G$ (which contain them). Note that by the construction of dual edges, every face of $G$ must have at least one vertex of $G^{*}$ contained in it. Consider a vertex $v_{i}^{*}$ in $G^{*}$, such that $m\notin\{i,\pi(i)\}$.
Now, we consider a decreasing greedy subsequence of the sequence $\pi(i),\pi^{2}(i),\dots,m$. We claim that this subsequence contains $f_{R}(i)$. Indeed, since $f_{R}(i)$ is the first element of $\pi(i),\pi^{2}(i),\dots,m$ that is less than $i$, it is less than all elements preceding it and therefore contained in a greedy decreasing subsequence. If we end this greedy subsequence before $f_{R}(i)$, it corresponds to a reverse-directed path from $\pi(i)_{out}$ to $i_{in}$. Similarly, an increasing greedy subsequence of the sequence $i,\pi^{-1}(i),\pi^{-2}(i),\dots,m$ contains $f_{L}(\pi(i))$, and we find a directed path from $i_{in}$ to $\pi(i)_{out}$. By construction, these two paths are edge-disjoint and therefore form a cycle. Furthermore, for the dual edge $j^{*}$ in $G^{*}$ of each edge $j$ on the first path we have $j_{out}=v_{i}^{*}$. Similarly, for the dual edge $j^{*}$ in $G^{*}$ of each edge $j$ on the second path we have $j_{in}=v_{i}^{*}$. Therefore, the vertex $v_{i}^{*}$ uniquely lies inside the face bounded by this cycle.
All that remains to be shown is that $\pi=\pi_{\mathcal{G}}$. By duality, we only need to consider any $i\mapsto\pi(i)$ such that $i>\pi(i)$. In $G$, the two endpoints of the edge $i$ are $i_{out}=v_{f_{R}(i)}=v_{\pi(i)}$ and $i_{in}=v_{\pi(f_{L}(i))}$. First, we claim that the edge $\pi(i)$ is the smallest edge with respect to $i$ that is incident to $i_{out}$. We that have $f_{L}(\pi(i))=i$ and therefore $\pi(i)_{in}=i_{out}$. For any other $j$ such that $j_{in}=i_{out}$, we must have $f_{L}(j)=i$, and therefore $\pi(i)<j<i$. For any $j$ such that $j_{out}=i_{out}$, we must have $f_{R}(j)=\pi(i)$, which also implies that $\pi(i)<j<i$. Therefore, $\pi_{M(G)}(i)\geq_{i}\pi(i)$. Consider the following cycle. Find a directed path from the edge $i$ to edge $m$ by taking the greedy increasing subsequence of $i,\pi^{-1}(i),\dots,m$. Then, find a reverse directed path from $\pi(i)$ to $m$ by taking the greedy decreasing subsequence of $\pi(i),\pi^{2}(i),\dots,m$. Clearly, the largest edge on this cycle with respect to $\leq_{i}$ is $\pi(i)$, and therefore $\pi_{\mathcal{G}}(i)=\pi(i)$. This completes the proof. ∎
883254671883254671
Figure 2. Graph $G$ (left) with $\pi(G)=(1,8,3,2,5,4,6,7)$, and $G^{*}$ (right, dashed) with $\pi(G^{*})=\pi^{-1}(G)$.
Rotate $\tau$ so that the first entry of the cycle is $n$. Note that this does not change the permutation. Label $\tau=(a_{1}=n,a_{2},\dots,a_{n})$ to obtain the sequence $a$, and construct $T(a)$. In order to construct $G$ from $T(a)$, we construct a path $P_{s}$ for each $s\in V(T(a))$ as follows. Let $P_{s}$ be the directed path of length $s$, where the edges are labelled with entries from $s$ in ascending order along the direction of the path. Note that this means that $s_{1}$ is the label of the last edge on $P_{s}$, adjacent to the sink vertex of $P_{s}$.
Theorem 3.2.
Let $\pi^{:}$ be a decorated permutation. Then, there exists a connected planar graph $G$ such that $\pi^{:}=\pi^{:}_{M(G)}$.
Proof.
We write the decorated permutation in disjoint cycle notation, $\pi^{:}=\pi^{:}_{1}\cdots\pi^{:}_{r}$, where for all integers $1\leq i\leq r$, $\pi^{:}_{i}$ is a cycle of length $k_{i}$. For each disjoint cycle $\pi^{:}_{i}$ on the set of elements $S_{i}$, we obtain the cycle $\tilde{\pi}^{:}_{i}$ on the set of elements $[k_{i}]$ by the unique, order preserving bijection $S_{i}\to[k_{i}]$. If $\tilde{\pi}^{:}_{i}$ is a type-$2$ fixed point, then $G_{i}=K_{2}$ and $E(G)=1$. Otherwise, we apply Lemma 3.1 to obtain $G_{i}$. We obtain the set of graphs $\{G_{i}\}^{r}_{i=1}$, such that $\tilde{\pi}^{:}_{i}=\tilde{\pi}^{:}_{M(G_{i})}$. For each graph $G_{i}$, we relabel $E(G_{i})$ using the bijection $[E(G_{i})]=[k_{i}]\to S_{i}$. Then, $\pi^{:}_{i}=\pi^{:}_{M(G_{i})}$. Furthermore, $\pi^{:}=\pi^{:}_{1}\cdots\pi^{:}_{r}=\pi^{:}_{M(G_{1})}\cdots\pi^{:}_{M(G_{r})}=\pi^{:}_{M(\bigsqcup G_{i})}$. Let $\{v_{i}\}^{r}_{i=1}$ be a set of vertices such that $v_{i}\in V(G_{i})$. Then, $G=\bigsqcup G_{i}/(v_{1}\sim\ldots\sim v_{r})$ is connected and Whitney $2$-isomorphic to $\bigsqcup G_{i}$, so $M(G)\cong M(\bigsqcup G_{i})$. Therefore, $\pi^{:}=\pi^{:}_{M(G)}$. ∎
Corollary 3.3.
Every open positroid variety contains a graphic matroid variety.
4. Binary, regular, and graphic positroids
In this section, we characterize the graphic positroids and show that the classes of binary, regular, graphic and series-parallel positroids are all equal.
A $2$-connected matroid is series-parallel if and only if it has no minor isomorphic to $U^{2}_{4}$ or $M(K_{4})$ [11]. As no positroid contains a minor isomorphic to $M(K_{4})$, we get the following result.
Corollary 4.1 (Corollary 6.4 in [16]).
A $2$-connected positroid is series-parallel if and only if it is $U^{2}_{4}$-free.
We may immediately strengthen this result, as any matroid that is $U^{2}_{4}$- and $M(K_{4})$-free is a direct sum of series-parallel matroids [11, Theorem 10.4.8].
Corollary 4.2.
A positroid is a direct sum of series-parallel matroids if and only if it is $U^{2}_{4}$-free.
We use the following result in Theorem 4.4.
Lemma 4.3.
Let $P$ be a $2$-connected graphic positroid on a densely totally ordered set $X$. Then, $\pi^{:}_{P}$ consists of a single cycle.
Proof.
As $P$ is a $2$-connected positroid, by Theorem 2.27, $P$ has a $2$-sum decomposition into positroids
$P=P_{1}\oplus_{2}(P_{2}\oplus_{2}\cdots(P_{n-1}\oplus_{2}P_{n})\cdots),$
such that for all $i\in[n-1]$, $E(N_{i})$ and $E(P_{i+1})$ are non-crossing subsets of $X$, where
$N_{i}=P_{1}\oplus_{2}(P_{2}\oplus_{2}\cdots(P_{i-1}\oplus_{2}P_{i})\cdots)$
is a positroid. Furthermore, as $P$ is graphic, for all $i\in[n]$, $P_{i}$ is a circuit or a cocircuit. We prove by induction that $\pi^{:}_{P}$ consists of a single cycle. Let $E(P_{1})=\{j_{1}<j_{2}<\cdots<j_{k}\}$. Suppose that $P_{1}$ is a circuit, then $\pi^{:}_{P_{1}}=(j_{k},j_{k-1},\ldots,j_{1})$. Suppose instead that $P_{1}$ is a cocircuit, then $\pi^{:}_{P_{1}}=(j_{1},j_{2},\ldots,j_{k})$. Now, suppose that for some $i\in[n-1]$, $\pi^{:}_{N_{i}}$ consists of a single cycle. Then, as $E(N_{i})$ and $E(P_{i+1})$ are non-crossing subsets of $X$, by Lemma 2.34, we have that $\pi^{:}_{N_{i}\oplus_{2}P_{i+1}}$ consists of a single cycle. It follows by induction that $\pi^{:}_{P}=\pi^{:}_{N_{n-1}\oplus_{2}P_{n}}$ consists of a single cycle. ∎
The following theorem summarizes the characterization of binary positroids and their envelope classes.
Theorem 4.4.
The following are equivalent for a positroid $P$.
  • (i)
    $P$ is binary,
  • (ii)
    $P$ is graphic,
  • (iii)
    $P$ is regular,
  • (iv)
    $P$ is a direct sum of series-parallel matroids,
  • (v)
    $P$ is the unique matroid contained in $\Omega_{P}$,
  • (vi)
    for every field $\mathbb{F}$, $\Pi^{\circ}_{\mathbb{F}}(P)=S_{P}(\mathbb{F})$,
  • (vii)
    $\Pi_{\mathbb{R}}(P)$ is equal to the disjoint union of graphic matroid strata,
Proof.
We break the proof up by items.
  • $(i)\Leftrightarrow(ii)$
    By Theorem 2.1, a binary positroid is $U^{2}_{4}$-free, thus by Corollary 4.2, the direct sum of series-parallel matroids and therefore graphic. By Theorem 2.2, a graphic positroid is $U^{2}_{4}$-free, hence binary.
  • $(i)\Leftrightarrow(iii)$
    Positroids are $\mathbb{R}$-realizable. By Theorem 6.6.3 in [11], a matroid is binary and $\mathbb{F}$-realizable for some field $\mathbb{F}$ of characteristic other than two if and only if $M$ is regular.
  • $(i)\Leftrightarrow(iv)$
    By Theorem 2.1, a binary positroid is $U^{2}_{4}$-free, hence by Corollary 4.2, the direct sum of series-parallel matroids. By Corollary 4.2, a positroid that is a direct sum of series-parallel matroids is $U^{2}_{4}$-free, hence by Theorem 2.1, binary.
  • $(i)\Leftrightarrow(v)$
    Suppose that $P$ is a binary and therefore graphic positroid on a densely totally ordered set $X$. Then by Theorem 2.19, $P$ is the direct sum of a collection of $2$-connected graphic positroids, $P=\bigoplus^{n}_{i=1}P_{i}$, that are on pairwise non-crossing, disjoint subsets of $X$. By Proposition 2.52, this induces a direct sum decomposition of the positroid envelope class of $P$, $\Omega_{P}=\bigoplus^{n}_{i=1}\Omega_{P_{i}}$. Thus, $|\Omega_{P}|=\Pi^{n}_{i=1}|\Omega_{P_{i}}|$. It is then sufficient to show that for all $i\in[n]$, $P_{i}$ is the unique matroid contained in $\Omega_{P_{i}}$.
    By Lemma 4.3, for all $i\in[n]$, $\pi^{:}_{P_{i}}$ consists of a single cycle, hence by Proposition 2.46, all matroids contained in $\Omega_{P_{i}}$ are $2$-connected. Let $M\in\Omega_{P_{i}}$. Then, as $P_{i}$ is binary, by Theorem 2.14 $M$ is binary, and by Theorem 2.15 $M=P_{i}$. Thus, $\Omega_{P_{i}}=P_{i}$. It follows that $P$ is the unique matroid contained in $\Omega_{P}$.
    Now, suppose that $P$ is the unique matroid contained in $\Omega_{P}$. Then by Theorem 3.2, $P$ is graphic, hence regular and therefore binary.
  • $(v)\Leftrightarrow(vi)$
    Let $P$ be a positroid of rank-$k$ on $n$ elements. Suppose that $P$ is the unique matroid contained in $\Omega_{P}$, hence regular and therefore realizable over every field. Then, for any field $\mathbb{F}$, we have
    $\Pi^{\circ}_{\mathbb{F}}(P)$ =$=\{V\in\operatorname{Gr}(k,\mathbb{F}^{n}):M_{V}\in\Omega_{P}\}$
    =$=\{V\in\operatorname{Gr}(k,\mathbb{F}^{n}):M_{V}=P\}=S_{P}(\mathbb{F}).$
    Suppose instead that for every field $\mathbb{F}$, $\Pi^{\circ}_{\mathbb{F}}(P)=S_{P}(\mathbb{F})$. Then, in particular, $\Pi^{\circ}_{\mathbb{F}}(P)=S_{P}(\mathbb{R})$. By Theorem 3.2, $P$ is graphic and therefore the unique matroid contained in $\Omega_{P}$.
  • $(vi)\Leftrightarrow(vii)$
    Let $P$ be a positroid of rank-$k$ on $X=\{i_{1}<i_{2}<\cdots<i_{n}\}$. Suppose that for every field $\mathbb{F}$, $\Pi^{\circ}_{\mathbb{F}}(P)=S_{P}(\mathbb{F})$. Then, in particular, $\Pi^{\circ}_{\mathbb{R}}=S_{P}(\mathbb{R})$. By Corollary 2.53,
    $\Pi_{\mathbb{R}}(P)=\{V\in\operatorname{Gr}(k,\mathbb{R}^{n}):I\notin\mathcal{I}(P)\Rightarrow\Delta_{I}(V)=0\}.$
    Observe that $\Pi_{\mathbb{R}}(P)$ can be equivalently defined as
[…]
    where the disjoint union is taken over all points $V\in\operatorname{Gr}(k,\mathbb{R}^{n})$ such that $\mathbb{1}:P\to M_{V}$ is a rank-preserving weak map. Let $V\in\operatorname{Gr}(k,\mathbb{R}^{n})$ be fixed, such that $\mathbb{1}:P\to M_{V}$ is a rank-preserving weak map. Let $\mathcal{J}(M_{V})=(J_{i_{1}},J_{i_{2}},\ldots,J_{i_{n}})$, and let $Q$ be the positroid envelope of $M_{V}$. Then, $\mathcal{J}(Q)=\mathcal{J}(M_{V})$. Let $\mathcal{J}(P)=(J^{\prime}_{i_{1}},J^{\prime}_{i_{2}},\ldots,J^{\prime}_{i_{n}})$, then for all $k\in[n]$, $J^{\prime}_{i_{k}}\leq_{i_{k}}J_{i_{k}}$. By Theorem 2.37, we obtain
    $\mathcal{B}(Q)$ $=\bigcap^{n}_{k=1}\left\{B={X\choose r}:J_{i_{k}}\leq_{i_{k}}B\right\}$
    $\subseteq\bigcap^{n}_{k=1}\left\{B={X\choose r}:J^{\prime}_{i_{k}}\leq_{i_{k}}B\right\}=\mathcal{B}(P).$
    Therefore, $\mathbb{1}:P\to Q$ is a rank-preserving weak map. As $P$ is the unique matroid contained in $\Omega_{P}$ and therefore binary, by Theorem 2.14, $Q$ is binary, hence graphic and therefore $M_{V}=Q$. It follows that $\Pi_{\mathbb{R}}(P)$ is the disjoint union of graphic matroid strata.
    Suppose instead that $\Pi_{\mathbb{R}}(P)$ is the disjoint union of graphic matroid strata. Then, $S_{P}(\mathbb{F})$ is a graphic matroid stratum, hence $P$ is a graphic and therefore the unique matroid contained in $\Omega_{P}$. It follows that for every field $\mathbb{F}$, $\Pi^{\circ}_{\mathbb{F}}(P)=S_{P}(\mathbb{F})$.
We may then immediately extend the following enumerative result for series-parallel positroids to graphic positroids.
Proposition 4.5.
Let $\beta$ be the map sending a permutation $w=w_{1}\ldots w_{n-1}$ in one-line notation to the permutation $\beta(w)=(w_{1},\ldots,n)$ in cycle notation. Then $\beta$ is a bijection between separable permutations on $[n-1]$ with $k$ descents and decorated permutations of series-parallel positroids in $\operatorname{Gr}^{\geq 0}(k+1,\mathbb{R}^{n})$.
Corollary 4.6.
Let $|X|=n$ be a totally ordered set. The connected graphic positroids on $X$ of rank $k+1$ are in bijection with the separable permutations on $[n-1]$ with $k$ descents.
For a totally ordered set $X$ of size $n$, it follows from Remark 3.24 in [7] and from [12] that the number of connected graphic positroids $X$ is given by the large Schrödner number $R_{n-2}$, and the refinement $R_{k,n-2}$ [8, A175124] counts the number of rank $k+1$ connected graphic positroids on $X$.
5. Incidence matrices
In this section, we strengthen the connection between the graphic positroids and the series-parallel graphs, by showing that every graphic positroid $P$ can be realized by a matrix with all nonnegative maximal minors that is the signed incidence matrix of a graph that represents $P$. Let $A$ be a full rank matrix and let $J$ be a length $r(A)$ sequence of distinct column vectors in $A$, then we denote by $\Delta_{J}(A)$ the maximal minor of $A$ given by $J$.
Observation 5.1.
Let $B$ be full rank, $m\times n$ matrix with column vectors $(c_{1},c_{2},\ldots,c_{n})$, where $n\geq m$. Suppose that $B$ has all nonnegative maximal minors. Then the full rank, $m\times n$ matrix $B^{\prime}$, with column vectors
$(c_{i+1},\ldots,c_{n},(-1)^{m-1}c_{1},\ldots,(-1)^{m-1}c_{i}),$
has all nonnegative maximal minors.
Proof.
It is sufficient to show the result for $i=1$. Let $J$ be a length $m$ sequence of distinct column vectors in $B^{\prime}$. Suppose that $J\subseteq\{c_{2},c_{3},\ldots,c_{n}\}$, then $J$ is also a sequence of distinct column vectors in $B$ and $\Delta_{J}(B^{\prime})=\Delta_{J}(B)\geq 0$. Suppose instead that
$J=(c_{j_{1}},\ldots,c_{j_{m-1}},(-1)^{m-1}c_{1})$
for some $\{c_{j_{t}}\}^{m-1}_{t=1}\subseteq\{c_{2},c_{3},\ldots,c_{n}\}$. Take $J^{\prime}=(c_{1},c_{j_{1}},\ldots,c_{j_{m-1}})$. Then $\Delta_{J}(B^{\prime})=(-1)^{m-1}\cdot(-1)^{m-1}\Delta_{J^{\prime}}(B)=\Delta_{J^{\prime}}(B)\geq 0$. ∎
Theorem 5.2.
Let $P$ be a graphic positroid of rank $r$. Then there exists a graph $G$ and matrix $A$, so that $M(G)=M(A)=P$, $A$ has all nonnegative maximal minors and $A$ is obtained from a signed incidence matrix of $G$ by deleting $|G|-r$ rows.
Proof.
It follows from Theorem 4.4 that $P$ can be represented by a graph $G$ that is the disjoint union of a set of series-parallel graphs. We show that there exists a full rank matrix $A$, obtained from a signed incidence matrix of $G$ by deleting rows, such that $M(A)=M(G)$ and $A$ has all nonnegative maximal minors. We begin by showing that the result holds for any series-parallel network. Any series-parallel network can be obtained from $C_{1}$, the cycle graph on a single vertex, or $K_{2}$, the complete graph on two vertices, by a sequence of series-parallel extensions. Let $G$ be a series-parallel network, and suppose that $G$ can be obtained from $K_{2}$ by a sequence of series-parallel extensions. Let $K_{2}=G_{1},G_{2},\ldots,G_{n}=G$ be a sequence of graphs such that for all $1\leq i<n$, $G_{i+1}$ is obtained from $G_{i}$ by a series or parallel extension of some edge $e\in E(G_{i})$. We inductively construct a sequence of matrices, $A_{G_{1}},A_{G_{1}},\ldots,A_{G_{n}}$, where $\forall i\in[n]$, $A_{G_{i}}$ is a signed incidence matrix for $G_{i}$. Furthermore, $\forall i\in[n]$, by deleting the second row we obtain a matrix $A_{i}$ that has all nonnegative maximal minors. For $G_{1}=K_{2}$, the matrix
$A_{G_{1}}=\left(\begin{array}[]{c}1\\ -1\end{array}\right)$
is a signed incidence matrix for $K_{2}$. Let $A_{1}$ be the matrix obtained from $A_{G_{1}}$ by deleting the second row. Then $A_{1}=(1)$, so it has it exactly one maximal minor, $\det(A)=1>0$.
Now consider the graph $G_{i+1}$ that is obtained from the graph $G_{i}$ by a series or parallel extension about some specified edge $e\in E(G_{i})$. Let us denote the column vectors of $A_{G_{i}}$, ordered from left to right, as the sequence $(c_{1},c_{2},\ldots,c_{i})$, and let $c_{e}$ be the column vector associated with the edge $e\in E(G_{i})$. Suppose that $G_{i+1}$ is obtained from $G_{i}$ by a parallel extension, that is $G_{i+1}$ is obtained from $G_{i}$ by adding an edge in parallel to $e$ in $G_{i}$. Suppose further that $A_{i}$, the matrix obtained from $A_{G_{i}}$ by deleting the second row, has all nonnegative maximal minors. We take $A_{G_{i+1}}$ to be the matrix whose column vectors, ordered from left to right, is the sequence $(c_{1},c_{2},\ldots,c_{e},c_{e},c_{e+1},\ldots,c_{i})$. Then $A_{G_{i+1}}$ is a signed incidence matrix of $G_{i+1}$. Furthermore, the set of non-zero maximal minors of $A_{i+1}$, the matrix obtained from $A_{G_{i+1}}$ by deleting the second row, is equal to the set of non-zero maximal minors of $A_{i}$. Therefore, all maximal minors of $A_{i+1}$ are nonnegative.
Suppose instead that $G_{i+1}$ is obtained from $G_{i}$ by a series extension, that is $G_{i+1}$ is obtained from $G_{i}$ by adding an edge in series to $e$ in $G_{i}$. Suppose further that $A_{i}$, the matrix obtained from $A_{G_{i}}$ by deleting the second row, has all nonnegative maximal minors. Let the column vector sequence of $A_{G_{i}}$, ordered from left to right, be given by $(c_{1},c_{2},\ldots,c_{i})$ and let $c_{e}$ be the column vector corresponding to the edge $e\in E(G_{i})$. The matrix $A^{\prime}_{G_{i+1}}$ given by the sequence of column vectors
$(c_{e+1},c_{e+2},\ldots,c_{i},(-1)^{r(A_{i})-1}c_{1},(-1)^{r(A_{i})-1}c_{2},\ldots,(-1)^{r(A_{i})-1}c_{e})$
is an incidence matrix of $G_{i}$. Furthermore, by Observation 5.1, the matrix $A^{\prime}_{i}$, obtained from $A^{\prime}_{G_{i}}$ by deleting the second row, has all nonnegative maximal minors. We construct a column vector $c^{\prime}_{e}$, where $\lvert c^{\prime}_{e}\rvert=\lvert c_{e}\rvert$, with entries given by
$c^{\prime}_{e}[j]:=\begin{cases}-1,&\text{if }(-1)^{r(A_{i})-1}c_{e}[j]=1\\ 0,&\text{otherwise.}\end{cases}$
Then the matrix
$A_{G_{i+1}}=\left(\begin{array}[]{@{}c|c|c@{}}A^{\prime}_{G_{i}}\setminus(-1)^{r(A_{i})-1}c_{e}&(-1)^{r(A_{i})-1}c_{e}+c^{\prime}_{e}&c^{\prime}_{e}\\ \hline\cr\begin{matrix}0\cdots 0\end{matrix}&1&1\end{array}\right)$
is a signed incidence matrix for $G_{i+1}$. Furthermore, the matrix $A^{\prime}_{i+1}$, obtained from $A_{G_{i+1}}$ by deleting the second row, has all nonnegative maximal minors. ∎
Acknowledgments
We thank Allen Knutson, for first posing the question of whether every positroid envelope class contains a graphic matroid, and for many helpful discussions. We also thank Melissa Sherman-Bennett for helpful comments.